### v1

# Computational Obfuscations and Random Oracles for Derandomizing Asynchronous Consensus

Abstract
A method for converting an asynchronous randomized consensus to a deterministic asynchronous consensus is presented. The method uses program computation obfuscation (e.g., [5, 12]) and a random oracle (e.g., [9, 6]), assuming a computationally bounded scheduler, to overcome the impossibility result of Fischer, Lynch, and Paterson [10].
Two stages are combined, in the first, a new form of computational program obfuscation implemented by post-quantum cryptographic hash functions (see e.g., [12]) is introduced, employing time lock puzzles (see e.g., [19, 1]) to imply a computational gap between the consensus participants and the (imaginary adversarial) scheduler. In the second stage, a random oracle is implemented by using a post-quantum cryptographic hash function that allows each process to harvest pseudo-randomization from the (cleartext) program and a (consensus) round (variable) and, in turn, implies the completion of the consensus in the presence of a computationally bounded scheduler.
1 Introduction
Asynchronous consensus is one of the most investigated tasks in distributed computing. Consensus over distributed inputs allows a reduction to centralized computation, abstracting away the inherent coordination difficulties distributed systems cope with.
Asynchronous consensus is a building block in structuring replicated state machines and efficient proof of authority/stake Blockchains, e.g., [8].
Unfortunately, the celebrated result of Fischer, Lynch, and Paterson [10] proves that there is no deterministic algorithm that can yield an asynchronous consensus. Paxos [14] succeeded in ensuring the consistency (all participants decide on the same value) and validity (the decided value appears in at least one of the inputs) of the distributed decision but sacrifices the liveness as an adversarial scheduler can keep the system undecided forever.
Randomization comes to the rescue (e.g., [7]). We investigate the possibility of using a random oracle [9] to derandomize a randomized asynchronous consensus using a post-quantum one-way function.
Structuring the random oracle led us to use post-quantum time lock puzzles (see, e.g., [19, 1] and the references therein) for efficiently structuring obfuscated programs (see, e.g., [12] and the references therein). Intuitively, to obfuscate the program, we computationally hide (by means of time lock puzzles) an important binary truth table of a threshold function, where the adversary scheduler needs to recover the threshold index, while the participants need only recover the value in two indices. The result is deterministic asynchronous consensus in the presence of a computationally bounded adversary.
Overview.
The system settings are described in the next section. The distributed consensus task is described in Section 3, including the impossibility result to achieve consistency on the decided (same) output (agreement), output that reflects at least one of the inputs (validity) and is chosen to be output in a finite number of steps (termination). Then, randomization is introduced, assisting in solving the challenging consensus problem. Section 5 describes a fundamental building block in randomized consensus, namely the conciliator. A conciliator is a Monte Carlo consensus object, based on approximate agreement, that guarantees validity and termination in all executions but guarantees agreement only with some probability. An implementation of a conciliator is described in Section 5. Obfuscation is used to yield a (tunable) computation logarithmic factor between the computation the scheduler has to invest (with relation to the participants) to avoid agreement in the conciliator output. An additional randomized consensus technique is used to overcome the unfortunate case where the conciliator outputs disagree. In this stage, the use of a random oracle forces the adversarial scheduler to invest more computation than each participant, alas risking a successful consensus. Thus, intuitively, the computation budget of the scheduler is exhausted in order to reach a consensus.
Implementations of the obfuscation and random oracle appear in Section 6 and Section 8, respectively.
2 The system
Processes
We define $p_{1},...,p_{n}$ to be a set of asynchronous processes. These processes communicate by reading and writing to shared objects.
Configuration
We define a configuration to be the state of the local memory of all $n$ processes $p_{1},\ldots,p_{n}$, including their program counters, together with the content of all shared objects.
Scheduler
We define a scheduler as an adversary that can decide which of the processes to activate next. This adversary can also follow the process program as dictated by the chosen process program counter and perform local computations that end with either a read operation from a shared object or a write operation to a shared object. The scheduler has complete information, namely, the scheduler knows the content of the shared objects, the local memory of each process, and the algorithm they execute.
Random Oracle
We define a random oracle $H$ to be an idealized version of the cryptographic hash function, that is, a black box that can be queried with input $x$ to return output $y$. The box is consistent; that is, any input $x$ will always output the same $y$. The box is uniformly random, which means the probability of getting output $y$ for an input $x$ that was not already queried is uniform. Everyone has access to this black box and can query it, but its internal workings are unknown (See Chapter 5 of [13]).
3 The consensus task
A consensus protocol is a protocol that is run by a set of $n$ asynchronous processes, each of which starts a binary input value $v_{i}\in\{0,1\}$, runs until it reaches a decision value $d_{i}\in\{0,1\}$, then halts. A consensus protocol is correct if it satisfies the following requirements:
Agreement
A consensus protocol satisfies agreement if no two processes decide on different values: $\forall i,j:d_{i}=d_{j}$.
Validity
A consensus protocol is valid if all processes decide on a value that was the input of some process, $\forall i\exists j:d_{i}=v_{j}$.
Termination
Every non-faulty process decides after a finite number of steps.
We think of the consensus problem as an adversarial game between the scheduler and the processes. The processes try to reach agreement on the same value by reading and writing to shared objects, and the scheduler tries to prevent them from doing so by strategically ordering these operations.
4 Bound on the cost of finding a bad execution
As defined by Fischer, Lynch, and Paterson [10], a configuration $C$ is bivalent if there is a path from it to a configuration where a process decides on the value $1$, and there is also a path from it to another configuration where a process decides on the value 00. Similarly, a configuration $C$ is univalent if all the paths from $C$ lead to configurations with an identical decision value. Because of the agreement property, any configuration in which some process has decided is necessarily univalent.
The FLP proof, in both its original form [10], and its extension to shared memory by Loui and Abu-Amara [15], gives an explicit procedure for constructing an infinite non-terminating execution for any protocol that satisfies agreement and validity, by starting in a bivalent configuration and always choosing bivalent successor configurations. However, the computational cost of this construction is not considered. We would like to find an adversary strategy that allows it to prevent a protocol from terminating without having to explore the (possibly infinite!) space of all executions.
The time complexity of computing a non-terminating, and thus infinite, execution is necessarily infinite. So, we will look at the more tractable problem of computing an extension for some fixed number of steps $s$. The goal is to replace the abstract adversary function with an adversary program that takes as input the starting configuration of the system and the programs used by the processes to choose their next operations, and, in the case of agreement, their outputs.
It is convenient to model this problem in terms of a weakened version of consensus where termination within a fixed number of steps and validity are always required, but agreement may fail. This is similar to the conciliator abstraction of [3], with the difference that we can no longer assign a probability to failure but must instead rely on failing executions being computationally difficult to find.
Formally, let us define the following problem. We will specify the behavior of each process using a random-access machine program $P$ that implements a function $Q×V\textrightarrow Q×A$, where $Q$ is the set of states of the process, $V$ is the set of possible return values of operations, and $A$ is the next operation to execute. An adversary is a program in the same model that takes as input the step bound $s$, a sequence of programs $\left\langle{P_{1},\dots,P_{n}}\right\rangle$ and initial states $q_{1},\dots,q_{n}$, and outputs a sequence of process ids specifying the order in which the processes take steps. We assume that each program $P_{i}$ has two special return actions $r_{0}$ and $r_{1}$, that it must choose between after exactly $s$ steps; these provide the output of each process in the consensus protocol. The adversary wins if any of these outputs disagree.
The following theorem shows that the cost to the adversary to find a bad execution, relative to the worst-case cost of each process to carry out some execution, scales linearly with the number of steps performed by each process.
Theorem 4.1.
Given inputs $P_{1},\dots,P_{n}$ representing processes in a shared-memory system where, in all executions, (a) each process $p_{i}$ outputs a decision value within $s$ steps, and (b) each process $p_{i}$ uses a total of at most $t$ time to compute its transitions, there is an adversary program that computes within $O(st)$ time a schedule that causes some pair of processes to output different decision values.
To prove the theorem, we consider a restricted version of the FLP bivalence argument in which valence is defined only in terms of solo extensions of the current configuration. This will significantly reduce the number of extensions the adversary needs to consider.
For each configuration $C$, define the preference $\textnormal{{pref}}_{p}(C)$ of $p$ in $C$ to be the output value of $p$ in the execution $𝐶𝛼$ where only $p$ takes steps before it decides. We will call such an extension $α$ $p$’s solo-terminating extension and write $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cα)$ to indicate that $p$ decides this value in configuration $𝐶𝛼$. We know that a unique such extension $α$ exists because $p$ is deterministic (making $α$ unique) and the system is wait-free (so that no fairness requirement prevents $p$ from running alone).
Call a configuration $C$ solo-bivalent if there are processes $p$ and $q$ such that $\textnormal{{pref}}_{p}(C)≠\textnormal{{pref}}_{q}(C)$. Call a configuration solo-$b$-valent if $\textnormal{{pref}}_{p}(C)=b$ for all $p$. Similar to the original FLP construction, our goal is to start in a solo-bivalent configuration and stay in a solo-bivalent configuration.
We start with some simple observations about how a process’s preference can change. Given executions $α$ and $β$, write $α\sim_{p}β$ ($α$ is indistinguishable by $p$ from $β$) if $α|p=β|p$, meaning that $p$ observes the same events in both executions. Note that $α\sim_{p}β$ in particular implies $\textnormal{{decision}}_{p}(α)=\textnormal{{decision}}_{p}(β)$.
Lemma 4.2.
Let $C$ be a configuration and let $π$ be an operation of $p$. Then $\textnormal{{pref}}_{p}(Cπ)=\textnormal{{pref}}_{p}(C)$.
Proof.
Let $α$ be $p$’s solo-terminating extension of $𝐶𝜋$. Then $πα$ is a $p$’s solo-terminating extension of $C$, and $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cπα)=\textnormal{{pref}}_{p}(Cπ)$. ∎
Lemma 4.3.
Let $π$ be a read operation by $p$. Then $\textnormal{{pref}}_{q}(Cπ)=\textnormal{{pref}}_{q}(C)$.
Proof.
Let $α$ be $q$’s solo-terminating extension of $C$. Then $Cπα\sim_{q}Cα$ implies $\textnormal{{pref}}_{q}(Cπ)=\textnormal{{decision}}_{q}(Cπα)=\textnormal{{decision}}_{q}(Cα)=\textnormal{{pref}}_{q}(C)$. ∎
Lemma 4.4.
Let $π_{p}$ and $π_{q}$ be writes by $p$ and $q$ to the same register. Then $\textnormal{{pref}}_{p}(Cπ_{q}π_{p})=\textnormal{{pref}}_{p}(C)$.
Proof.
Since the value written by $π_{q}$ is replaced by $π_{p}$, we have $Cπ_{q}π_{p}\sim_{p}Cπ_{p}$ and thus $\textnormal{{pref}}_{p}(Cπ_{q}π_{p})=\textnormal{{pref}}_{p}(Cπ_{p})=\textnormal{{pref}}_{p}(C)$. ∎
Lemma 4.5.
For any wait-free shared-memory protocol satisfying termination and validity, there exists an initial solo-bivalent configuration for any $n≥2$.
Proof.
Take any configuration $C$ where two processes $p$ and $q$ have different inputs. If either process runs alone, it observes only its own input and is forced to decide it by validity. ∎
Lemma 4.6.
Let $C$ be a solo-bivalent configuration where $p$ and $q$ are processes with $\textnormal{{pref}}_{p}(C)≠\textnormal{{pref}}_{q}(C)$. Let $π_{p}$ and $π_{q}$ be the pending operators of $p$ and $q$ in $C$. Then, at least one of the following holds:
  1. 1.
    $Cπ_{p}$ is solo-bivalent with $\textnormal{{pref}}_{p}(Cπ_{p})=\textnormal{{pref}}_{p}(C)$ and $\textnormal{{pref}}_{q}(Cπ_{p})=\textnormal{{pref}}_{q}(C)$.
  2. 2.
    $Cπ_{q}$ is solo-bivalent with $\textnormal{{pref}}_{p}(Cπ_{q})=\textnormal{{pref}}_{p}(C)$ and $\textnormal{{pref}}_{q}(Cπ_{q})=\textnormal{{pref}}_{q}(C)$.
  3. 3.
    $Cπ_{p}π_{q}$ is solo-bivalent with $\textnormal{{pref}}_{p}(Cπ_{p}π_{q})≠\textnormal{{pref}}_{p}(C)$ and $\textnormal{{pref}}_{q}(Cπ_{p}π_{q})≠\textnormal{{pref}}_{q}(C)$.
Proof.
By Lemma 4.3, $\textnormal{{pref}}_{p}(Cπ_{p})=\textnormal{{pref}}(C)$, so to avoid case (3), we must have $\textnormal{{pref}}_{q}(Cπ_{p})≠\textnormal{{pref}}_{q}(C)$.
Similarly, if case (2) does not hold, we must have $\textnormal{{pref}}_{p}(Cπ_{q})≠\textnormal{{pref}}_{q}(C)$.
If neither of these cases holds, then we have both
$\textnormal{{pref}}_{p}(Cπ_{q})$ $\textnormal{{pref}}_{p}(C)$
and
$\textnormal{{pref}}_{q}(Cπ_{p})$ $\textnormal{{pref}}_{q}(C).$
From Lemma 4.3, neither $π_{p}$ nor $π_{q}$ is a read. From Lemma 4.4, $π_{p}$ and $π_{q}$ cannot be writes to the same location. So $π_{p}$ and $π_{q}$ are writes to different locations. But then $Cπ_{p}π_{q}$ and $Cπ_{q}π_{p}$ yield the same configuration. It follows that
$\textnormal{{pref}}_{p}(Cπ_{p}π_{q})$ $=\textnormal{{pref}}_{p}(Cπ_{q}π_{p})=\textnormal{{pref}}_{p}(Cπ_{q})≠\textnormal{{pref}}_{p}(C)$
while
$\textnormal{{pref}}_{q}(Cπ_{p}π_{q})$ $=\textnormal{{pref}}_{q}(Cπ_{p})≠\textnormal{{pref}}_{q}(C).$
So case (3) holds. ∎
One difference between solo-bivalence and bivalence that is illustrated by the third case of Lemma 4.6 is that it is not necessarily the case that a configuration with a solo-bivalent successor is itself solo-bivalent. But we are happy as long as we can find solo-bivalent extensions of solo-bivalent configurations, even if this means passing through intermediate configurations that may not be solo-bivalent.
To turn Lemma 4.6 into an algorithm, we just need to be able to test which of its three cases holds. To compute a bad schedule, the adversary carries out the following steps:
  1. 1.
    Pick two processes $p$ and $q$, and start in a configuration $C$ where $\textnormal{{pref}}_{p}(C)=0$ and $\textnormal{{pref}}_{q}(C)=1$, as in Lemma 4.5.
  2. 2.
    Given $C$ with $\textnormal{{pref}}_{p}(C)≠\textnormal{{pref}}_{q}(C)$:
    1. (a)
      If $\textnormal{{pref}}_{q}(Cπ_{p})=\textnormal{{pref}}_{q}(C)$, append $π_{p}$ to the schedule and set $C\textleftarrow Cπ_{p}$.
    2. (b)
      If $\textnormal{{pref}}_{p}(Cπ_{q})=\textnormal{{pref}}_{p}(C)$, append $π_{q}$ to the schedule and set $C\textleftarrow Cπ_{q}$.
    3. (c)
      Otherwise, the third case of Lemma 4.6 holds. Append $π_{p}π_{q}$ to the schedule and set $C\textleftarrow Cπ_{p}π_{q}$.
  3. 3.
    Repeat until one of $p$ and $q$ decides; then run the other to completion, appending its operations to the schedule.
The cost of computing $\textnormal{{pref}}_{p}(C)$ for any configuration $C$ is $O(t)$, since we can just simulate $p$ until it decides. To avoid copying, we do this simulation in place, logging any changes to $p$’s state or the state of the shared memory so that we can undo them in $O(t)$ time. For computing $\textnormal{{pref}}_{p}(Cπ_{q})$, we again pay at most $O(t)$ time, since applying $π_{q}$ to $C$ takes no more that $O(t)$ time, as does running $p$ to completion thereafter. The same bounds hold with the roles of $p$ and $q$ reversed.
It follows that each iteration of the main body of the algorithm finishes in $O(t)$ time. Since each of $p$ and $q$ do at most $s$ steps before deciding, there are at most $2s=O(s)$ iterations, giving a total cost of $O(st)$, even taking into account the $O(t)$ of the last phase of the algorithm. Since every iteration of the algorithm yields a solo-bivalent configuration, the preferences of $p$ and $q$ are never equal, and so the constructed schedule violates agreement. This concludes the proof of Theorem 4.1.
5 Hidden threshold conciliator
The ratio between the $O(t)$ time complexity of each process in Theorem 4.1 and the $O(st)$ time complexity of the adversary suggests that a useful strategy for the processes faced with a computationally-limited adversary is to exhaust the adversary’s resources by forcing it to simulate a large number of expensive potential executions.
In this section, we describe how to implement a derandomized conciliator for $n$ processes that reduces the problem of achieving agreement on the existence of a cryptographic implementations of a threshold function where the threshold can only be identified by probing specific function values, which we construct in Section 6. The cost to the adversary to find a bad execution of this conciliator is similar to the upper bound shown in Theorem 4.1.
shared data: array $A[1\dots s][2]$ of atomic registers, initialized to
[…]
2       $i\textleftarrow\textnormal{{input}}$
3       for $r\textleftarrow 1\dots s$ do
             // write my current position
4             $A[r][i\bmod 2]\textleftarrow i$
             // read opposite position
5             $i^{\prime}\textleftarrow A[r][(i+1)\bmod 2]$
6             if $i^{\prime}=⊥$ then
                   // no opposition, stay put
7                   $i\textleftarrow 2⋅i$
8
9            else
                   // adopt midpoint
10                   $i\textleftarrow i+i^{\prime}$
11
12
13      return $i$
14
[…]
16       $i\textleftarrow\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
17       return $f(i)$
18
Algorithm 5.1 Derandomized threshold conciliator
The pseudocode for this conciliator is given in Algorithm 5.1. The main procedure conciliator takes as arguments a consensus input value $\textnormal{{input}}∈\left\{{0,1}\right\}$, a representation of a hidden threshold function $f$, and a security parameter $s$ that scales the number of shared-memory steps taken by the process.
The conciliator proceeds in two phases.
First, the processes carry out a $s$-round approximate agreement protocol that assigns each process $p$ a value $v_{p}$ in the range $0\dots 2^{s}$, with the properties that
  1. 1.
    Any process that sees only input $v$ obtains value $2^{s}⋅v$.
  2. 2.
    For any two processes $p$ and $q$, $\left\lvert v_{p}-v_{q}\right\rvert≤1$.
The approximate agreement protocol is an adaptation of Moran’s one-dimensional midpoint protocol [16]. In Moran’s original protocol, each process takes a sequence of snapshots of current values and adopts the midpoint of the values it sees until all fit within a particular range. In our protocol, we organize this sequence into a layered execution of $s$ rounds, where in each round, a process adopts the average value it sees, scaled by a factor of $2$ per round to track the integer numerators instead of the actual fractional values. By taking advantage of the inputs starting at 00 and $1$, we can show that at most two values, differing by at most $1$, appear at the end of each round; this allows us to replace the snapshot with a pair of registers, one of which holds the even value and one the odd. Formally:
Lemma 5.1.
Let $S_{r}$ be the set of all $i$ values that are held by any process after $r$ iterations of the loop in procedure approximateAgreement. Then $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$ for some $v_{r}$.
Proof.
By induction on $r$. For $r=0$, we have $S_{0}⊆\left\{{0,1}\right\}$.
For $r+1$, from the induction hypothesis, there is a value $v_{r}$ such that $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$. So, each process writes either $v_{r}$ or $v_{r+1}$ to one of the registers $A[r][-]$ before reading the other register; since $v_{r}$ and $v_{r}+1$ have different values mod $2$, they will not overwrite each other.
It follows that whichever value $v_{r}$ or $v_{r+1}$ is written first will be visible to all processes. If this is $v_{r}$, then every process sees either $v_{r}$ alone and chooses a new value $2⋅v_{r}$), or sees $v_{r}$ and $v_{r+1}$ and chooses $v_{r}+(v_{r}+1)=2⋅v_{r}+1$. In this case, we get $S_{r+1}⊆\left\{{v_{r+1},v_{r+1}+1}\right\}$ where $v_{r+1}=2⋅v_{r}$. The case where $v_{r}+1$ is written first is similar, except now $v_{r+1}=2⋅v_{r}+1$ instead. ∎
In particular, the set of output values is given by $S_{s}⊆\left\{{v_{s},v_{s}+1}\right\}$ for some $v_{s}$.
To obtain its decision value, each process $p$ then feeds its output $v_{p}$ to a function $f(v_{p})$. To ensure validity, we require that $f(0)=0$ and $f(2^{s})=1$. Agreement is obtained if $f(v_{p})=f(v_{q})$ for all processes $p$ and $q$.
This makes the task of the adversary simple: given $f$, it must find consecutive positions $v$ and $v+1$ such that $f(v)≠f(v+1)$, then steer the execution of the approximate agreement protocol to these values to prevent agreement. We aim to design $f$ to make this task computationally expensive relative to the costs paid by the processes. We assume that the primary cost to both the adversary and each process will be evaluating $f$ for particular inputs. Since each process evaluates $f$ exactly once, the main question becomes how many times the adversary evaluates $f$.
The worst case is when $f$ is a threshold function with an unknown threshold. Here, the adversary’s best strategy to find a bad pair $f(v)≠f(v+1)$ is a binary search, which requires $Θ(s)$ evaluations of $f$.
Lemma 5.2.
Let $T$ be a threshold drawn uniformly from $\left\{{1,\dots,2^{s}}\right\}$, and let Let $f(v)=1$ if and only if $v≥T$. Then an adversary with oracle access to $f$ requires $Ω(s)$ evaluations of $f$ on average to find $v$ such that $f(v)≠f(v+1)$, while an adversary that is restricted to $s/2$ evaluations of $f$ finds such a $v$ with probability at most $2^{-s/2}$.
Proof.
Because $f$ is chosen without regard to the adversary’s strategy, we can limit our consideration to deterministic adversaries and then apply Yao’s Principle [20] to generalize to probabilistic adversaries. The rest of the argument is essentially the usual lower bound for binary search.
Fix some deterministic adversary strategy. We can express this strategy as a decision tree where each node specifies some value $v$ at which to compute $f(v)$, with two children corresponding to the result of observing $f(v)=0$ or $f(v)=1$. Each leaf of the tree is labeled by an output value $v$.
Observe that the adversary’s strategy succeeds in finding $v$ such that $f(v)≠f(v+1)$ precisely when $v=T-1$. So to succeed in all executions, the decision tree must have at least $2^{s}$ leaves to cover all possible values of $T$, which implies an average depth of at least $s$ by Kraft’s inequality. This gives the $Ω(s)$ lower bound to find a bad pair always.
If the tree has depth $s/2$, then it has at most $2^{s/2}$ leaves. The chance that a random $T$ lands on one of the at most $2^{s/2}$ values corresponding to one of these leaves is at most $2^{-s/2}$. ∎
The lemma immediately gives bounds on the cost of computing a bad schedule for Algorithm 5.1:
Corollary 5.3.
Assume that $f$ satisfies the conditions of Lemma 5.2, and let $t$ be a lower bound on the expected time complexity to compute $f(v)$. Then it costs $Ω(st)$ on average to find an execution of Algorithm 5.1 that violates agreement, and there is a constant $c>0$ such that any adversary limited to $cst$ total time finds an execution that violates agreement with probability at most $2^{-s/2}$.
In the following section, we show how to construct a function $f$ that satisfies the conditions of Lemma 5.2 and Corollary 5.3.
6 Computational obfuscation for threshold encoding
Because we cannot restrict the adversary from observing the process’s programs, we cannot directly enforce that the adversary only accesses $f$ as an oracle. So instead, we must rely on an obfuscated implementation of $f$ that is both expensive to evaluate and reveals no information about $f(v^{\prime})$ for $v^{\prime}=v$ when evaluating $v$ other than that implied by the threshold-function property.
Specifically, we introduce a new approach for computationally obfuscating programs that use one-way function primitives. The program’s creation uses randomization that can be revealed by (tunable inversion processing required to) invert the one-way function. We further demonstrate the usefulness of such an obfuscation by enabling the execution of the obfuscated program, somewhat analogously to how programs operate on fully homomorphic encrypted data.
We start with the case in which the inversion of the one-way function can be easily performed, namely encrypting an array of values that is used to encode a threshold value $T$ (an array index) such that all entries that are less than $T$ encode 00 while the others encode $1$. Any asymmetric or symmetric encryption scheme (e.g., RSA, ElGamal, AES) can encrypt random strings that encode 00, say have bit parity 00, to be the first $T-1$ entries of the threshold encoding array, and the other entries to encode $1$. Then, provide the threshold encoding array and the decryption key as part of the computationally obfuscated program. The benefit of such obfuscation is more apparent when the computation needed to compute the obfuscated program is significantly smaller than the computation the adversary (in our case, the scheduler) needs to invest. Thus, we suggest using one-way functions that are employed in the context of proof of work, namely, SHA.
The $T-1$ first entries of the threshold encoding array are based on a random string $\langle choice,nonce,r,ed\rangle$, where choice is 00, while the the remaining $\ell-T+1$ entries have choice $1$. The value of $nonce$ is a random string chosen independently at random for each entry and appears in plaintext as an entry of a $nonce$ array in the program; the number of bits in the $nonce$ tunes the inversion difficulty. Using $nonce$ ensures that one (possibly exhaustive) search for an inverse is not helpful for an entry with a different $nonce$. $r$ is a random string, and $ed$ is an error detection code over the concatenated $nonce$ and $r$. $ed$ is used to avoid unplanned collisions of SHA that may lead to the existence of more than one $T$ value. A long enough $nonce$ can suffice to fulfill the $ed$ mission and allow to eliminate the $ed$ and use only $\langle choice,nonce,r\rangle$. In addition, $choice$ can be defined by the xor of all bits of $nonce$ and $r$, choosing a $nonce$ to fit the required 00 values in the $T-1$ prefix entries and $1$ values in the rest. In such a case, each entry of the threshold encoding array results from SHA applied to the concatenation of only the $nonce$ and $r$ associated with the entry.
Formally, let $H_{m}:\{0,1\}^{m}\rightarrow\{0,1\}^{m}$ be a secure hash function for some fixed $m\in\mathbb{N}$. In particular, we discuss the NIST post-quantum standard hash [18] SHA 512/256. Let $0<T≤\ell$ be a threshold value and let $f_{T}:\mathbb{Z}_{\ell}\rightarrow\{0,1\}$ be our threshold function for $T$, that is
$f_{T}(i)=\begin{cases}0&i<T\\ 1&i≥T\end{cases}$
1 procedure $\textnormal{{preprocess}}(\ell,T,m,k$ s.t. $t\leq\ell\wedge k<m)$
       // Initialize array of hashes and array of nonces
2       $C[1,...,\ell][1,...m]$ an uninitizalized array of bit strings
3       $P[1,...,\ell][1,...k]$ an uninitizalized array of bit strings
4       for $i\textleftarrow 1\dots\ell$ do
5             $P[i]\leftarrow$ random bit string of length $k$
6             $r_{i}\leftarrow$ random bit string of length $m-k$
7
8            if $i≥T$ then
9                   $a_{i}\leftarrow 1$
10
11            else
12                   $a_{i}\leftarrow 0$
13
             // enforce that XOR of $r_{i}$ is equal to $f_{t}(i)$
14             if $\bigoplus_{j=1}^{m-k}[r_{i}]_{j}\neq a_{i}$ then
15                   flip a bit in $r_{i}$ at random
16
17            $y_{i}\leftarrow P[i]\|r_{i}$
18             $C[i]\leftarrow H_{m}(y_{i})$
19
20      return $C,P$
21
Algorithm 6.1 Threshold Encapsulation
Preprocessing Algorithm.
Algorithm 6.1 is executed by the programmer to obtain a computational obfuscated protocol using preprocessing. The integer parameters $\ell$, $T$, $k$, and $m$ are declared (Line 1). Where $\ell$ is the span in which a threshold $t$ is defined, $m$ is the number of bits used by the one-way function, say 256 for SHA 256, and $k$ is the length of the (exposed) prefix nonce of the $m$ bits input (and output) of the SHA.
Two arrays, $C$ (committed values) and $P$ (padding nonces), each of $\ell$ entries, are defined in lines 6.1 and 6.1. Each entry of the $C$ array is ready to be assigned with the SHA result (line 6.1). Each entry of the $P$ array exposes a $k$-bit nonce value (line 6.1). We use the nonce to enforce the independence of pre-image searches and as randomness used to validate the pre-image.
The for loop (lines 6.1 to 6.1) iterates over all possible values of $i$ and computes the cryptographic hash (e.g., the common one-way function such as SHA) over the nonce concatenated with $r_{i}$, a concealed random string of length $m-k$. The computation starts (line 6.1) with a random selection of bits string of size $k$ as the nonce for $i$ and entry and stores it in $P[i]$. Next, the algorithm randomly selects a bit string $r_{i}$ (line 6.1); this string must have the property that $\oplus_{j}[r_{i}]_{j}=1$ if and only if $i≥T$, and thus the algorithm sets $a_{i}=f(i)$ accordingly (lines 6.1 to 6.1), and checks (line 6.1) if the randomly chosen $r_{i}$ satisfies this property. If it does not, it simply flips one bit at random in $r_{i}$ (line 6.1). Since $\oplus_{j}[r_{i}]_{j}\neq a_{i}$, flipping any bit $[r_{i}]_{l}\leftarrow\neq[r_{i}]_{l}$ will cause the xor of the bits in $r_{i}$ to change and thus we get $\oplus_{j}[r_{i}]_{j}=a_{i}$.
The algorithm concatenates the nonce and $r_{i}$ to obtain a bit string $y_{i}$ of length $m$ (line 6.1), then the algorithm computes the hash of $y_{i}$ using $H_{m}$ and stores the resulting hash in $C[i]$ (line 6.1). By the time the algorithm reaches the return command (line 6.1), it has computed the hashes for all the $n$ hash values of $C$, and all the nonces used to calculate them are stored in $P$; thus, the algorithm returns them to be embedded in the protocol program.
local data: integers $\ell,m,k$, precomputed array of hashes $C[1,...,\ell][1,...,m]$ and nonces $P[1,...,\ell][1,...,k]$
[…]
2       $A[1,...,m]$ an uninitialized bit string
3       $A\leftarrow$ pre-image of hash $C[i]$ of size $m$ s.t., $A[1,...,k]=P[i][1,...,k]$
4 return $\bigoplus_{j=k}^{m}A[j]$
Algorithm 6.2 Threshold Probe
Execution of the Obfuscated Program.
The processes use the probingThreshold procedure described in Algorithm 6.2 to compute $f_{T}(i)$.
The algorithm works by finding the pre-image $P[i]\|r_{i}$ that was used to compute $C[i]$ in the preprocessing phase. It begins by initializing a bit string $A$ of length $m$ to hold potential pre-images to $C[i]$ (line 2). Next, it finds and stores in $A$ a pre-image for $C[i]$ such that the first $k$ bits of the pre-image $A$ are equal to $P[i]$ (line 6.2) since if $A=P[i]\|r_{i}$ then $A[1,...,k]=P[i][1,...,k]$. This means that, with high probability $A=P[i]\|r_{i}$ and therefore $A[k,...,m]=r_{i}$. The exact method of finding this pre-image has not been specified, but exhaustive search mining or any other method to find a pre-image of a cryptographic hash function can be employed. Thus, Algorithm 6.2 returns $a_{i}$ by computing $\bigoplus_{j=k}^{m}A[j]$ (line 6.2), since the preprocessing algorithm defined $a_{i}=f_{T}(i)$, the algorithm returns $f_{T}(i)$.
Security Arguments Sketch.
Consider series of hashes $C=C_{1},\dots,C_{\ell}$ and $P=P_{1},\dots,P_{\ell}$ and nonces generated by Algorithm 6.1 for some $\ell,T,m,k$. We call the tuple $\mathcal{S}=(\ell,m,k,C,P)$ an obfuscation of the function $f_{T}$. We define a query to $\mathcal{S}$ with input $i\in[n]$ as finding the pre-image $P[i]\|r_{i}$ of a hash $C_{i}\in C$ and computing $\bigoplus_{j}[r_{i}]_{j}$.
Consider a malicious computationally bounded adversary that implements an adaptive adversary scheduler commonly assumed in the scope of asynchronous distributed computing. The scheduler computation bound is used to prove the security of the obfuscated construction. If the adversary can reveal $T$ in an expected number of $o(\log\ell)$ pre-images computations, then one can implement a search better than the binary search, which is proven optimal, a contradiction that implies the need to compute expected $\Theta(\log\ell)$ pre-images.
Furthermore, if the adversary can find a pre-image earlier than planned by (the SHA based time lock puzzle), then the one-way function used, say, SHA, is easy to reverse, which violates the common standard and belief.
7 From conciliators to consensus
In Section 5, we gave an implementation of a deterministic conciliator-like protocol where each process carries out $O(s)$ steps and a computation with $O(t)$ time complexity, while the adversary is forced to spend $Ω(st)$ time to find a bad execution that prevents agreement. This implementation requires significant setup, and because each instance is fixed and deterministic, once the adversary finds a bad execution, it costs nothing to reuse this bad execution if we are foolish enough to reuse the conciliator. We are also faced with the issue that we really want a protocol that guarantees agreement in all executions.
We can address both of these issues by supplementing the conciliator with a chain of adopt-commit and conciliator objects as suggested in [3]. An adopt-commit object [17, 2, 4] implements the adopt-commit protocol of Gafni [11], in which each process arrives with an input value $v$ and leaves with an output that is either $\left\langle{\textnormal{{adopt}},v^{\prime}}\right\rangle$ or $\left\langle{\textnormal{{commit}},v^{\prime}}\right\rangle$ for some value $v^{\prime}$. It satisfies the same validity and wait-freedom conditions as consensus, but agreement is replaced with two new conditions: consistency, which requires that if any process returns $\left\langle{\textnormal{{commit}},v}\right\rangle$, all other processes return $\left\langle{\textnormal{{commit}},v}\right\rangle$ or $\left\langle{\textnormal{{adopt}},v}\right\rangle$ for the same value $v$; and coherence, which requires that if all processes have the same input $v$, they all return $\left\langle{\textnormal{{commit}},v}\right\rangle$.
When used in alternation with conciliators, adopt-commit objects allow processes to decide immediately on any value that they see in a $\left\langle{\textnormal{{commit}},v}\right\rangle$, since consistency implies that all other processes will at least adopt the same value and the validity of the conciliators will preserve this agreement until the next adopt-commit, which will cause any remaining processes to also return $v$ because of coherence. So the only issue left is how to build conciliators that will push the processes into agreement.
For the first round, we use the threshold conciliator from Section 5. If this fails to produce agreement, in subsequent rounds we switch to a simpler random-oracle-based conciliator, in which a computationally-expensive hash function $H$ is applied to a nonce and the current round number to generate a value that acts like a shared coin agreed on by all the processes.
1procedure $\textnormal{{consensus}}(\textnormal{{input}},f,s,\textnormal{{nonce}})$
       // Attempt to reach agreement with thresholdConciliator
2       $v\textleftarrow\textnormal{{thresholdConciliator}}(\textnormal{{input}},f,s)$
3       $(a,v)\textleftarrow\textnormal{{adoptCommit}}_{1}(v)$
4       if $a=\textnormal{{commit}}$ then
5             return $v$
6
      // Switch to multi-round backup protocol
7       for $r\textleftarrow 2\dots∞$ do
             // Attempt to reach agreement with oracleConciliator
8             $v\textleftarrow\textnormal{{oracleConciliator}}(r,v,\textnormal{{nonce}})$
9             $(a,v)\textleftarrow\textnormal{{adoptCommit}}_{r}(v)$
10             if $a=\textnormal{{commit}}$ then
11                   return $v$
12
13
14
[…]
16       $\textnormal{{present}}[r][v]\textleftarrow 1$
17       if $\textnormal{{present}}[r][¬v]=0$ then
             // no disagreement, keep current preference
18             return $v$
19      else
             // disagreement, use random oracle
20             return $H(r,\textnormal{{nonce}})$
21
Algorithm 7.1 Derandomized consensus protocol
Pseudocode for this procedure is given in Algorithm 7.1. The main procedure consensus assumes an infinite array of shared adopt-commit objects $\textnormal{{adoptCommit}}_{1},\textnormal{{adoptCommit}}_{2},\dots$, where $\textnormal{{adoptCommit}}_{r}$ tests for agreement in round $r$. The input to the consensus procedure is the usual consensus input input, plus the hidden-threshold function $f$ and security parameter $s$ for thresholdConciliator plus a nonce value used to prevent replay attacks on oracleConciliator. The oracleConciliator procedure assumes access to a random oracle that takes as input the current round number and a nonce value, and returns 00 or $1$ with equal probability.
Theorem 7.1.
Let $t$ be the cost of computing $f$ and $t^{\prime}$ the cost of computing $H$. Then Algorithm 7.1 satisfies agreement and validity in all executions, and there is a constant $c>0$ such that if the adversary is limited to $cst$ time complexity, Algorithm 7.1 terminates in expected $O\left(s+2^{-s/2}st/t^{\prime}\right)$ steps and $O\left(t+2^{-s/2}st\right)$. time for each process.
If the time complexity of the adversary is bounded by a larger value $b$, then Algorithm 7.1 still terminates in expected $O(b/t^{\prime})$ steps and $O(b)$ time for each process.
Proof.
Agreement and validity follow immediately from the properties of the adopt-commit objects and the fact that each conciliator satisfies validity. (For oracleConciliator, this is because if all processes have the same input, all process see $\textnormal{{present}}[r][¬v]=0$ in line 7.1 and return this input.)
If the adversary is limited to at most $cst$ time complexity, then from Corollary 5.3, the algorithm finishes after the first round with probability at least $1-2^{-s/2}$, incurring a cost of $O(s)$ steps and $O(t)$ time to each process. In the event that it does not finish, the adversary loses its ability to compute $H(r,\textnormal{{nonce}})$ after $O(cst/t^{\prime})$ rounds, after which each round is the last with probability $1/2$, since there is at most one value $v$ for which the processes with value $v$ skip the oracle, and the oracle output equals $v$ with this probability. It follows that in this case, the processes consume an additional $O(cst/t^{\prime})$ steps and $O(cst)$ time. Multiplying by the bound on the probability of this case gives the expected step complexity for each process of $O\left(s+2^{-s/2}st/t^{\prime}\right)$ and expected time complexity of $O\left(t+2^{-s/2}st\right)$.
If the adversary is given a sufficiently large time bound $b$, it can instead allocate $O(st)$ time to force the round-$1$ conciliator to fail. Predicting the output of $H$ in each subsequent round requires $t^{\prime}$ time, equal up to small additive constant to the time cost to each process that executes $H$. But after $O(b/t^{\prime})$ rounds, the adversary again runs out of time, and so the protocol terminates with an expected cost of $O(b/t^{\prime})$ steps and $O(b)$ time for each process. ∎
If the algorithm designer can tune $s$ based on a known bound $b$ on the time complexity of the adversary, the natural choice is to choose $s$ and $t$ so that the first case of the theorem holds adversary pays $Θ(st)$ time (assuming $2^{-s/2}s=O(1)$) while each process pays only $O(t)$.
8 Possible implementations of random oracles
Let nonce be a (tunable) part of $P[1,\ldots,\ell]$, for a given round $r$ consider $F=SHA(r,\textnormal{{nonce}})$, where the number of bits in $F$ is significantly smaller than the number of bits of the chosen nonce, allowing (by the pigeon hole principle) a set of preimages $PI$ (that includes $r,\textnormal{{nonce}}$) to yield the same value of $F$. Define $H(r,\textnormal{{nonce}})$ to be the parity of the (lexicographically) smallest preimage in $PI$.
The length of the portion used from, say, $P[1,\ldots,\ell]$ (the $C$ array, or the common program code, can serve us too) may tune the exhaustive search for the smallest preimage, as an exhaustive search over longer potential preimages may require more computation resources. Another obvious computation tuning capability is the number of output bits of the cryptographic hash function, e.g., SHA128 where the output is 128 bits, or one that outputs a different number of bits.
Processes use this computationally demanding procedure to harvest (pseudo) random bits, forcing the scheduler to invest more computation to coordinate a bivalent preserving strategy.
9 Concluding remarks
Designing an algorithm to cope with an adversarial scheduler (and/or inputs) ensures the algorithm’s function in rare scenarios where the schedule is the most unfortunate for the algorithm. In many cases, the worst case leads to an impossibility result, while the rare scenario does not happen in practice. This is particularly true in scenarios that require ongoing tracing and computation of the imaginary scheduler entity.
We examine the power of using program obfuscations and random oracles to derandomize distributed asynchronous consensus algorithms. The power of the random oracle can be demonstrated in algorithms where symmetry is broken by other means, such as process identifiers and/or symmetry-breaking operations, such as compare-and-swap. Thus, randomization has an inherent more substantial power in breaking symmetry. On the other hand, harvesting proper randomization during runtime is challenging, and it should be avoided if possible.
A weakness of our construction is that it requires substantial preprocessing to construct an obfuscated threshold function since our implementation is just a truth table hidden behind time-lock puzzles. A natural question is whether a more sophisticated obfuscation procedure could reduce this setup, perhaps by constructing a function $f(s,i)$ where the threshold $T$ is a hash of some shared data $s$ but the implementation of $f$ prevents recovering $T$ more efficiently than simply doing binary search.
However, even with these limitations, we believe that we have demonstrated integrating program obfuscation and random oracle abstractions and functionalities in distributed computing is helpful for the theory and practice of distributed computing and systems.
Acknowledgment: It is a pleasure to thank Sergio Rajsbaum for the initial interactions.