### v2

# Obfuscated Consensus

Abstract
The classic Fischer, Lynch, and Paterson impossibility proof FischerLP1985 ; LouiA1987 demonstrates that any deterministic protocol for consensus in either a message-passing or shared-memory system must violate at least one of termination, validity, or agreement in some execution. But it does not provide an efficient procedure to find such a bad execution.
We show that for wait-free shared memory consensus, given a protocol in which each process performs at most $s$ steps computed with total time complexity at most $t$, there exists an adversary algorithm that takes the process’s programs as input and computes within $O(st)$ time a schedule that violates agreement. We argue that this bound is tight assuming the random oracle hypothesis: there exists a deterministic obfuscated consensus protocol that forces the adversary to spend $Ω(st)$ time to find a bad execution despite having full access to all information available to the protocol.
This bound is based on a general algorithm that reduces the constructing an obfuscated consensus protocol to constructing an obfuscated threshold function that provably costs $Ω(t)$ time to evaluate on a single input, where $t$ is a tunable parameter, and for which an adversary with access to the threshold function implementation cannot extract the threshold any faster than by doing binary search. We give a particular implementation of such an obfuscated threshold function that is not very efficient but that is provably secure assuming the random oracle hypothesis. Since our obfuscated consensus protocol does not depend on the specific details of this construction, it may be possible to replace it with one that is more efficient or requires weaker cryptographic assumptions, a task we leave for future work.
1 Introduction
Asynchronous consensus is one of the most investigated tasks in distributed computing. Consensus over distributed inputs allows a reduction to centralized computation, abstracting away the inherent coordination difficulties distributed systems cope with.
Asynchronous consensus is a building block in structuring replicated state machines and efficient proof of authority/stake Blockchains, e.g., DBLP:journals/tdsc/DolevGNW24 .
Unfortunately, the celebrated result of Fischer, Lynch, and Paterson FischerLP1985 proves that there is no deterministic algorithm that can yield an asynchronous consensus. Subsequent consensus protocols have escaped this bound by adjusting the requirements of the problem. Paxos DBLP:journals/tocs/Lamport98 succeeded in ensuring the consistency (all participants decide on the same value) and validity (the decided value appears in at least one of the inputs) of the distributed decision but sacrifices the liveness as an adversarial scheduler can keep the system undecided forever. Alternatively, randomization (for example, DBLP:conf/podc/Ben-Or83 ) allows consensus to be solved at the cost of replacing a deterministic termination guarantee with a probabilistic one, where the protocol finishes with probability $1$ in the limit. In these protocols, the adversary scheduler that could otherwise prevent agreement following the procedure given by the FLP proof is unable to predict the future and thus falls off the bad path.
We investigate an alternative approach, where validity and termination are guaranteed in all executions, and agreement holds for all but a small number of executions that are computational expensive to compute, even for an adversary given access to the entire code and knowledge of the processes in the protocol. This idea is inspired by the Fiat-Shamir heuristic FS for extracting a zero-knowledge challenge from a hash of public values assuming a random oracle. In our approach, the scheduling choices of the adversary in an asynchronous shared-memory system, as observed by the processes, become the input to an obfuscated threshold function. Each process computes its output by computing a single value of this function, while the adversary can only prevent agreement through the much more expensive process of finding the threshold input at which the value of the function changes from 0 to $1$. This gap can in principle be exploited to obtain a fully-determinsitic wait-free asynchronous protocol that solves consensus in most executions assuming the adversary has limited computational power.
This approach has some limitations. We show that, given any protocol in which each process performs $s$ steps and computes for $t$ time, there is an adversary program that outputs an agreement-violating schedule in $O(st)$ time in the worst case.
Conversely, we provide a reduction from our obfuscated consensus problem to constructing an obfuscated threshold function, and give a particular implementation of such a function as an obfuscated truth table. Here each bit of the truth table is hidden using a post-quantum time lock puzzle (see, e.g., rivest1996time ; DBLP:conf/asiacrypt/AfsharCHLM23 and the references therein) for efficiently structuring obfuscated programs (see, e.g., DBLP:journals/cacm/JainLS24 and the references therein). This allows us to tune the base cost of computing the threshold function for both the processes and the adversary, and to prevent the adversary from finding the hidden threshold any faster than by binary search over a range exponential in the step complexity of the consensus protocol.
While our particular implementation of an obfuscated threshold function involves a relatively expensive setup phase, it is possible to imagine an implementation using more sophisticated cryptographic techniques that would avoid this overhead. We discuss this possibility further in Section˜7.
1.1 Overview
The system settings are described in the next section. The problem of obfuscated consensus is described in Section 3: this replaces the usual agreement condition for consensus that requires all processes output the same value in each execution with a new obfuscated agreement condition that only requires that it is computationally expensive for the adversary scheduler to find a schedule that violates agreement. Section 5 gives a reduction from obfuscated consensus to a purely cryptographic problem, that of obfuscating a threshold function. This reduction has parameters allowing both the computation cost incurred by each process and the ratio of the adversary’s cost to each process’s cost to be adjusted independently. An implementation of an obfuscated threshold function, along with a proof of security, is given in Section 6. Finally, we discuss some of the limitations of our current implementation and possibilities for future in Section 7.
2 Model
We use a standard asynchronous shared-memory model, with some small tweaks to include explicitly the programs used by the processes and adversary to compute their choices.
The system consists of $n$ asynchronous processes $p_{1},...,p_{n}$ that communicate by reading and writing to shared objects, with timing controlled by an adversary scheduler.
The behavior of each process is controlled by a random-access machine program $P$ that implements a function $Q×V\textrightarrow Q×A$, where $Q$ is the set of states of the process, $V$ is the set of possible return values of operations (including a special null value used to indicate the first step by the process), and $A$ is the next operation to execute.
The adversary scheduler is a program in the same model that takes as input a step bound $s$, a sequence of programs $\left\langle{P_{1},\dots,P_{n}}\right\rangle$ and initial states $q_{1},\dots,q_{n}$, and outputs a a sequence of process ids, where each id occurs at most $s$ times, specifying the order in which the processes take steps. We will refer to this sequence of process ids as a schedule.
Communication is via shared memory: A collection of atomic read/write registers $r_{1},\dots,r_{m}$, where applying a read operation to a register returns its current value and applying a write operations sets the value and returns nothing.
A configuration of the system consists of the states of the processes and registers, together with at most one pending operation for each process.
Every schedule yields an execution, an alternating sequence of configurations and operations $C_{0}α_{1},C_{1}α_{2}C_{2}\dots$ in which each $C_{i+1}$ is the result of applying operation $α_{i+1}$ to configuration $C_{i}$. Because the processes are deterministic, the adversary can simulate the execution corresponding to any particular schedule in time proportional to the total time used by the processes to choose their steps in that execution.
We assume that both the processes and the adversary have access to a random oracle $H$, which is an idealized version of the cryptographic hash function. The random oracle acts as a black box that can be queried with input $x$ to return output $y$. The box is consistent; that is, any input $x$ will always output the same $y$. The box is uniformly random, which means the probability of getting output $y$ for an input $x$ that was not already queried is uniform. Though everyone has access to this black box and can query it, its internal workings are unknown (See Chapter 5 of KatzLindell ).
3 Obfuscated consensus
A consensus protocol is a protocol that is run by a set of $n$ asynchronous processes, each of which starts a binary input value $v_{i}\in\{0,1\}$, runs until it reaches a decision value $d_{i}\in\{0,1\}$, then halts. A consensus protocol is correct if it satisfies the following requirements:
Agreement
A consensus protocol satisfies agreement if no two processes decide on different values: $\forall i,j:d_{i}=d_{j}$.
Validity
A consensus protocol is valid if all processes decide on a value that was the input of some process, $\forall i\exists j:d_{i}=v_{j}$.
Termination
Every non-faulty process decides after a finite number of steps.
We think of the consensus problem as an adversarial game between the scheduler and the processes. The processes try to reach agreement on the same value by reading and writing to shared objects, and the scheduler tries to prevent them from doing so by strategically ordering these operations.
The Fischer-Lynch-Paterson (FLP) impossibility result FischerLP1985 , as extended to shared memory by Loui and Abu-Amara LouiA1987 , shows that in an asynchronous system with n2n≥2 processes in which at least one process can fail by crashing, there is always an adversary strategy that finds an execution in which at least one of the three conditions is violated. We will be interested in constructing protocols where finding this strategy is computationally difficult.
Define an obfuscated consensus prototocol to be a protocol that satisfies termination and validity always, but replaces agreement with a condition that we can describe informally as
Obfuscated agreement
The cost to the adversary to compute a schedule that violates agreement substantially exceeds the cost to the processes to execute the protocol.
More formally, the adversary must solve an obfuscated consensus problem: Given process programs $P_{1},\dots,P_{n}$ and a step bound $s$, find a schedule in which each process takes $s$ steps that violates one of termination, validity, or agreement. A successful obfuscated consensus protocol is one for which solving this problem is expensive, for some reasonable definition of expensive. We will start by bounding this cost from above.
4 Bounding the cost of finding a bad execution
The core idea of the FLP impossibility proof FischerLP1985 is the notion of bivalence. A configuration $C$ is bivalent if there is a path from it to a configuration where a process decides on the value $1$, and there is also a path from it to another configuration where a process decides on the value 0. Similarly, a configuration $C$ is univalent if all the paths from $C$ lead to configurations with an identical decision value. Because of the agreement property, any configuration in which some process has decided is necessarily univalent.
The FLP proof, in both its original form FischerLP1985 , and its extension to shared memory by Loui and Abu-Amara LouiA1987 , gives an explicit procedure for constructing an infinite non-terminating execution for any protocol that satisfies agreement and validity, by starting in a bivalent configuration and always choosing bivalent successor configurations. However, the computational cost of this construction is not considered. We would like to find an adversary strategy that allows it to prevent a protocol from terminating without having to explore the space of all executions.
The time complexity of computing a non-terminating, and thus infinite, execution is necessarily infinite. We will instead look at the more tractable problem of computing an extension for some fixed number of steps $s$, which is the justification for defining obfuscated consensus as satisfying termination and validity always but possibly failing agreement. As described in Section 2, we replace an abstract adversary function with a concrete adversary program that takes as input the starting configuration of the system and the programs used by the processes to choose their next operations and, ultimately, their outputs.
Formally, let us define the following problem. We assume that each program $P_{i}$ has two special return actions $r_{0}$ and $r_{1}$, that it must choose between after exactly $s$ steps; these provide the output of each process in the consensus protocol. The adversary, a program that takes as input the programs $P_{1},\dots,P_{n}$ and the step bound $s$ and constructs a schedule in which each process takes exactly $s$ steps, wins if any of the processes’ outputs disagree.
The following theorem shows that the cost to the adversary to find a bad execution, relative to the worst-case cost of each process to carry out some execution, scales linearly with the number of steps performed by each process.
Theorem 4.1.
Given inputs $P_{1},\dots,P_{n}$ representing processes in a shared-memory system where, in all executions, (a) each process $p_{i}$ outputs a decision value in $s$ steps, and (b) each process $p_{i}$ uses a total of at most $t$ time to compute its transitions, there is an adversary program that computes in $O(st)$ time a schedule that causes some pair of processes to output different decision values.
To prove the theorem, we consider a restricted version of the FLP bivalence argument in which valence is defined only in terms of solo extensions of the current configuration. This will significantly reduce the number of extensions the adversary needs to consider.
For each configuration $C$, define the preference $\textnormal{{pref}}_{p}(C)$ of $p$ in $C$ to be the output value of $p$ in the execution Cα where only $p$ takes steps before it decides. We will call such an extension $α$ $p$’s solo-terminating extension and write $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cα)$ to indicate that $p$ decides this value in configuration Cα. We know that a unique such extension $α$ exists because $p$ is deterministic (making $α$ unique) and the system is wait-free (so that no fairness requirement prevents $p$ from running alone).
Call a configuration $C$ solo-bivalent if there are processes $p$ and $q$ such that $\textnormal{{pref}}_{p}(C)≠\textnormal{{pref}}_{q}(C)$. Call a configuration solo-$b$-valent if $\textnormal{{pref}}_{p}(C)=b$ for all $p$. Similar to the original FLP construction, our goal is to start in a solo-bivalent configuration and stay in a solo-bivalent configuration.
We start with some simple observations about how a process’s preference can change. Given executions $α$ and $β$, write $α\sim_{p}β$ ($α$ is indistinguishable by $p$ from $β$) if $α|p=β|p$, meaning that $p$ observes the same events in both executions. Note that $α\sim_{p}β$ in particular implies $\textnormal{{decision}}_{p}(α)=\textnormal{{decision}}_{p}(β)$.
Lemma 4.2.
Let $C$ be a configuration and let $π$ be an operation of $p$. Then $\textnormal{{pref}}_{p}(Cπ)=\textnormal{{pref}}_{p}(C)$.
Proof.
Let $α$ be $p$’s solo-terminating extension of Cπ. Then $πα$ is a $p$’s solo-terminating extension of $C$, and $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cπα)=\textnormal{{pref}}_{p}(Cπ)$. ∎
Lemma 4.3.
Let $π$ be a read operation by $p$. Then $\textnormal{{pref}}_{q}(Cπ)=\textnormal{{pref}}_{q}(C)$.
Proof.
Let $α$ be $q$’s solo-terminating extension of $C$. Then $Cπα\sim_{q}Cα$ implies $\textnormal{{pref}}_{q}(Cπ)=\textnormal{{decision}}_{q}(Cπα)=\textnormal{{decision}}_{q}(Cα)=\textnormal{{pref}}_{q}(C)$. ∎
Lemma 4.4.
Let $π_{p}$ and $π_{q}$ be writes by $p$ and $q$ to the same register. Then $\textnormal{{pref}}_{p}(Cπ_{q}π_{p})=\textnormal{{pref}}_{p}(C)$.
Proof.
Since the value written by $π_{q}$ is replaced by $π_{p}$, we have $Cπ_{q}π_{p}\sim_{p}Cπ_{p}$ and thus $\textnormal{{pref}}_{p}(Cπ_{q}π_{p})=\textnormal{{pref}}_{p}(Cπ_{p})=\textnormal{{pref}}_{p}(C)$. ∎
Lemma 4.5.
For any wait-free shared-memory protocol satisfying termination and validity, there exists an initial solo-bivalent configuration for any n2n≥2.
Proof.
Take any configuration $C$ where two processes $p$ and $q$ have different inputs. If either process runs alone, it observes only its own input and is forced to decide it by validity. ∎
Lemma 4.6.
Let $C$ be a solo-bivalent configuration where $p$ and $q$ are processes with $\textnormal{{pref}}_{p}(C)≠\textnormal{{pref}}_{q}(C)$. Let $π_{p}$ and $π_{q}$ be the pending operators of $p$ and $q$ in $C$. Then, at least one of the following holds:
  1. 1.
    $Cπ_{p}$ is solo-bivalent with $\textnormal{{pref}}_{p}(Cπ_{p})=\textnormal{{pref}}_{p}(C)$ and $\textnormal{{pref}}_{q}(Cπ_{p})=\textnormal{{pref}}_{q}(C)$.
  2. 2.
    $Cπ_{q}$ is solo-bivalent with $\textnormal{{pref}}_{p}(Cπ_{q})=\textnormal{{pref}}_{p}(C)$ and $\textnormal{{pref}}_{q}(Cπ_{q})=\textnormal{{pref}}_{q}(C)$.
  3. 3.
    $Cπ_{p}π_{q}$ is solo-bivalent with $\textnormal{{pref}}_{p}(Cπ_{p}π_{q})≠\textnormal{{pref}}_{p}(C)$ and $\textnormal{{pref}}_{q}(Cπ_{p}π_{q})≠\textnormal{{pref}}_{q}(C)$.
Proof.
By Lemma 4.3, $\textnormal{{pref}}_{p}(Cπ_{p})=\textnormal{{pref}}(C)$, so to avoid case (3), we must have $\textnormal{{pref}}_{q}(Cπ_{p})≠\textnormal{{pref}}_{q}(C)$.
Similarly, if case (2) does not hold, we must have $\textnormal{{pref}}_{p}(Cπ_{q})≠\textnormal{{pref}}_{q}(C)$.
If neither of these cases holds, then we have both
$\textnormal{{pref}}_{p}(Cπ_{q})$ $≠\textnormal{{pref}}_{p}(C)$
and
$\textnormal{{pref}}_{q}(Cπ_{p})$ $≠\textnormal{{pref}}_{q}(C).$
From Lemma 4.3, neither $π_{p}$ nor $π_{q}$ is a read. From Lemma 4.4, $π_{p}$ and $π_{q}$ cannot be writes to the same location. So $π_{p}$ and $π_{q}$ are writes to different locations. But then $Cπ_{p}π_{q}$ and $Cπ_{q}π_{p}$ yield the same configuration. It follows that
$\textnormal{{pref}}_{p}(Cπ_{p}π_{q})$ $=\textnormal{{pref}}_{p}(Cπ_{q}π_{p})=\textnormal{{pref}}_{p}(Cπ_{q})≠\textnormal{{pref}}_{p}(C)$
while
$\textnormal{{pref}}_{q}(Cπ_{p}π_{q})$ $=\textnormal{{pref}}_{q}(Cπ_{p})≠\textnormal{{pref}}_{q}(C).$
So case (3) holds. ∎
One difference between solo-bivalence and bivalence that is illustrated by the third case of Lemma 4.6 is that it is not necessarily the case that a configuration with a solo-bivalent successor is itself solo-bivalent. But we are happy as long as we can find solo-bivalent extensions of solo-bivalent configurations, even if this means passing through intermediate configurations that may not be solo-bivalent.
To turn Lemma 4.6 into an algorithm, we just need to be able to test which of its three cases holds. To compute a bad schedule, the adversary carries out the following steps:
  1. 1.
    Pick two processes $p$ and $q$, and start in a configuration $C$ where $\textnormal{{pref}}_{p}(C)=0$ and $\textnormal{{pref}}_{q}(C)=1$, as in Lemma 4.5.
  2. 2.
    Given $C$ with $\textnormal{{pref}}_{p}(C)≠\textnormal{{pref}}_{q}(C)$:
    1. (a)
      If $\textnormal{{pref}}_{q}(Cπ_{p})=\textnormal{{pref}}_{q}(C)$, append $π_{p}$ to the schedule and set $C\textleftarrow Cπ_{p}$.
    2. (b)
      If $\textnormal{{pref}}_{p}(Cπ_{q})=\textnormal{{pref}}_{p}(C)$, append $π_{q}$ to the schedule and set $C\textleftarrow Cπ_{q}$.
    3. (c)
      Otherwise, the third case of Lemma 4.6 holds. Append $π_{p}π_{q}$ to the schedule and set $C\textleftarrow Cπ_{p}π_{q}$.
  3. 3.
    Repeat until one of $p$ and $q$ decides; then run the other to completion, appending its operations to the schedule.
The cost of computing $\textnormal{{pref}}_{p}(C)$ for any configuration $C$ is $O(t)$, since we can just simulate $p$ until it decides. To avoid copying, we do this simulation in place, logging any changes to $p$’s state or the state of the shared memory so that we can undo them in $O(t)$ time. For computing $\textnormal{{pref}}_{p}(Cπ_{q})$, we again pay at most $O(t)$ time, since applying $π_{q}$ to $C$ takes no more that $O(t)$ time, as does running $p$ to completion thereafter. The same bounds hold with the roles of $p$ and $q$ reversed.
It follows that each iteration of the main body of the algorithm finishes in $O(t)$ time. Since each of $p$ and $q$ do at most $s$ steps before deciding, there are at most $2s=O(s)$ iterations, giving a total cost of $O(st)$, even taking into account the $O(t)$ of the last phase of the algorithm. Since every iteration of the algorithm yields a solo-bivalent configuration, the preferences of $p$ and $q$ are never equal, and so the constructed schedule violates agreement. This concludes the proof of Theorem 4.1.
5 Obfuscated consensus from obfuscated threshold
The ratio between the $O(t)$ time complexity of each process in Theorem 4.1 and the $O(st)$ time complexity of the adversary suggests that a useful strategy for the processes faced with a computationally-limited adversary is to exhaust the adversary’s resources by forcing it to simulate a large number of expensive potential executions. We can do this by making the outcome of each process’s execution depend on evaluating an expensive function of the view that process has in the execution.
An obfuscated threshold function $f:\left\{{0,\dots,\ell}\right\}\textrightarrow\left\{{0,1}\right\}$ is a function, represented as a random-access machine program, with $f(0)=0$ and $f(\ell)=1$. The obfuscated threshold problem is to find, given a representation of such a function, an input $v$ such that f(v)f(v+1)f(v)≠f(v+1). We will show that finding a bad execution of a particular length of any obfuscated consensus protocol is equivalent to solving this problem.
Binary search will find a $v$ with f(v)f(v+1)f(v)≠f(v+1) in $O(\log\ell)$ evaluations of $f$, which puts an upper bound on the cost of solving the obfuscated threshold function problem. But it may be that a sufficiently clever algorithm can extract $v$ from $f$ without evaluating $f$ explicitly. We will discuss how to build an $f$ that resists such attacks in Section 6.3.
Note that the end conditions $f(0)=0$ and $f(1)=1$ show that at least one transitional value $v$ always exists, and, for some choices of $f$, many such transitional values exist. This makes the name “obfuscated threshold function” a bit misleading, since $f$ is not necessarily a threshold function. But choosing a threshold function in particular reduces the likelihood of guessing a bad $v$ at random, so we will try to produce threshold functions if we can.
We will show a reduction between these problems by constructing an obfuscated consensus protocol that uses an obfuscated threshold function to compute its output values. In Algorithm 5.1, we give a protocol that assumes that every process (and thus also the adversary) has access to an obfuscated threshold function $f$, and show that the adversary computes a bad execution for this protocol if and only if it can find a $v$ such that f(v)f(v+1)f(v)≠f(v+1).
shared data: array $A[1\dots s][2]$ of atomic registers, initialized to
1 procedure $\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
2 $i\textleftarrow\textnormal{{input}}$
3 for $r\textleftarrow 1\dots s$ do
    // write my current position
4    $A[r][i\bmod 2]\textleftarrow i$
    // read opposite position
5    $i^{\prime}\textleftarrow A[r][(i+1)\bmod 2]$
6    if $i^{\prime}=$⊥ then
       // no opposition, stay put
7       $i\textleftarrow 2⋅i$
8
9    else
       // adopt midpoint
10       $i\textleftarrow i+i^{\prime}$
11
12
13 return $i$
14
15procedure $\textnormal{{obfuscatedConsensus}}(\textnormal{{input}},f,s)$
16 $i\textleftarrow\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
17 return $f(i)$
18
Algorithm 5.1 Obfuscated consensus based on an obfuscated threshold
Pseudocode for a protocol that does this is given in Algorithm 5.1. The main procedure obfuscatedConsensus takes as arguments a consensus input value $\textnormal{{input}}∈\left\{{0,1}\right\}$, a representation of a function $f$ that converts integer values to Boolean decisions, and a security parameter $s$ that scales the number of shared-memory steps taken by the process.
The protocol proceeds in two phases.
First, the processes carry out a $s$-round approximate agreement protocol that assigns each process $p$ a value $v_{p}$ in the range $0\dots 2^{s}$, with the properties that
  1. 1.
    Any process that sees only input $v$ obtains value $2^{s}⋅v$.
  2. 2.
    For any two processes $p$ and $q$, $\left\lvert v_{p}-v_{q}\right\rvert≤1$.
The approximate agreement protocol is an adaptation of Moran’s one-dimensional midpoint protocol Moran1995 . In Moran’s original protocol, each process takes a sequence of snapshots of current values and adopts the midpoint of the values it sees until all fit within a particular range. In our protocol, we organize this sequence into a layered execution of $s$ rounds, where in each round, a process adopts the average value it sees, scaled by a factor of $2$ per round to track the integer numerators instead of the actual fractional values. By taking advantage of the inputs starting at 0 and $1$, we can show that at most two values, differing by at most $1$, appear at the end of each round; this allows us to replace the snapshot with a pair of registers, one of which holds the even value and one the odd. Formally:
Lemma 5.1.
Let $S_{r}$ be the set of all $i$ values that are held by any process after $r$ iterations of the loop in procedure approximateAgreement. Then $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$ for some $0≤v_{r}≤2^{r}-1$.
Proof.
By induction on $r$. For $r=0$, we have $v_{0}=0$ and $S_{0}⊆\left\{{0,1}\right\}$.
For $r+1$, from the induction hypothesis, there is a value $v_{r}$ such that $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$. So, each process writes either $v_{r}$ or $v_{r+1}$ to one of the registers A[r][]A[r][-] before reading the other register; since $v_{r}$ and $v_{r}+1$ have different values mod $2$, they will not overwrite each other.
It follows that whichever value $v_{r}$ or $v_{r+1}$ is written first will be visible to all processes. If this is $v_{r}$, then every process either sees $v_{r}$ alone and chooses a new value $2⋅v_{r}$, or sees $v_{r}$ and $v_{r+1}$ and chooses $v_{r}+(v_{r}+1)=2⋅v_{r}+1$. In this case, we get $S_{r+1}⊆\left\{{v_{r+1},v_{r+1}+1}\right\}$ where $v_{r+1}=2⋅v_{r}$. The case where $v_{r}+1$ is written first is similar, except now the processes all choose either $v_{r+1}=2⋅v_{r}+1$ or $v_{r+1}+1=2⋅(v_{r}+1)$. ∎
In particular, the set of output values is given by $S_{s}⊆\left\{{v_{s},v_{s}+1}\right\}$ where $0≤v_{s}≤2^{s}-1$.
To obtain its decision value, each process $p$ then feeds its output $v_{p}$ to a function $f(v_{p})$. To ensure validity, we require that $f(0)=0$ and $f(2^{s})=1$. Agreement is obtained if $f(v_{p})=f(v_{q})$ for all processes $p$ and $q$.
If $f$ requires $T(f)$ time to evaluate, the time complexity incurred by each process running Algorithm 5.1 is $O(s)+T$. From Theorem 4.1 this implies that there is an adversary strategy that computes a bad execution in time $O(s^{2}+sT)$; this also tells us that at least one bad execution exists. We’d like to show a comparable lower bound on the cost to the adversary, assuming it is difficult to find a threshold in $f$.
Since every process has a value in $\left\{{v_{s},v_{s}+1}\right\}$, agreement holds automatically if $f(v_{s})=f(v_{s}+1)$. So an adversary that finds an execution that produces disagreement also finds a value $v_{s}$ with $f(v_{s})≠f(v_{s}+1)$.
For any fixed $f$ with $f(0)=0$ and $f(2^{s})=1$, there is a random-access machine that outputs a threshold value $v$ with f(v)f(v+1)f(v)≠f(v+1) in $O(1)$ time, since it can just include $v$ in its code. We will need to avoid this by considering the average-case cost to find a threshold value for an $f$ drawn from some family of threshold functions. Applying Algorithm 5.1 to the functions in this family then give a family of consensus protocols.
Theorem 5.2.
Let $F$ be a family of threshold functions with range $\left\{{0,\dots,\ell}\right\}$. Fix some some $n$ and some $s$ such that $2^{s}$$\ell$, and for each fFf∈F, let $P_{f}$ be the obfuscated consensus protocol given by instantiating the $s$-round version of Algorithm 5.1 with $f$ as its decision function.
Suppose we sample $P_{f}$ uniformly at random and consider a starting configuration in which not all processes have the same input. If there is a probabilistic random-access machine $M$ that takes $P_{f}$ and the process’s initial states as input, runs for $T$ expected time, and outputs a schedule that produces disagreement, then there is a probabilistic random access machine $M^{\prime}$ that takes $f$ as input, runs for $T+O(ns)$ expected time, and outputs a $v$ with f(v)f(v)+1f(v)≠f(v)+1.
Proof.
We give a construction for $M^{\prime}$.
Run $M$ on $P_{f}$ to obtain a schedule for Algorithm 5.1, using $T$ time. Simulate the algorithm until each process $p$ returns a value $i_{p}$ from approximateAgreement, but do not have any process call $f$. this takes $O(ns)$ time since each simulated step can be done in constant time in the RAM model. Output the smallest $i_{p}$ value as $v$.
The cost of this procedure is bounded by $T+O(ns)$. From Lemma 5.1 we know that all processes obtain either $v$ or $v+1$, and in the case that $M$ outputs a schedule that produces disagreement, we have f(v)f(v+1)f(v)≠f(v+1). ∎
In the following section, we give an implementation of a family of threshold functions with these properties.
6 Implementing an obfuscated threshold function
Because we cannot restrict the adversary from observing the process’s programs, we cannot directly enforce that the adversary only accesses $f$ as an oracle. So instead, we must rely on an obfuscated implementation of $f$ that is both expensive to evaluate and reveals no information about $f(v^{\prime})$ for $v^{\prime}=v$ when evaluating $v$ other than that implied by the threshold-function property.
6.1 Definition
We introduce a new approach for computationally obfuscating programs that use one-way function primitives. The program’s creation uses randomisation that can be revealed by (tunable) inversion processing required to invert the one-way function. We construct the obfuscation by using a random oracle to hide a sequence of bits that encode the truth table of a threshold function. To obtain the output from the obfuscation, one must find the pre-image of a single hash. Thus it is imperative that the hardness of computing the pre-image be tunable such that it is feasible to find a single pre-image, yet infeasible for the computationally bounded adversary to enough pre-images to de-obfuscate the program. We achieve this tunability not by weakening the hash function, but by providing a prefix for the pre-image, such that the processes need only complete it. We begin with the following definitions.
Definition 6.1 (Threshold Function).
Let $T,\ell\in\mathbb{N}$ be two integers such that $0<T<\ell$. Define function $f_{T}:[\ell]\rightarrow\{0,1\}$ as
$f_{T}(i)=\begin{cases}0&i<T\\ 1&i\geq T\end{cases}$
Then $f_{T}$ is a threshold function for threshold $T$.
Definition 6.2 (Random Oracle Hash Function).
Let $m\in\mathbb{N}$ and let $H_{m}:\{0,1\}^{m}\rightarrow\{0,1\}^{m}$ be sampled uniformly at random from the set of all functions $\{0,1\}^{m}\to\{0,1\}^{m}$ (i.e., $H_{m}$ is a random oracle / random function). Then for all distinct $x\neq x^{\prime}\in\{0,1\}^{m}$ and all $y,y^{\prime}\in\{0,1\}^{m}$,
  1. 1.
    $\Pr\left[H_{m}(x)=y\right]=2^{-m}$
  2. 2.
    $\Pr\left[H_{m}(x)=y\ \wedge\ H_{m}(x^{\prime})=y^{\prime}\right]=2^{-2m}$
Definition˜6.2 is adapted from the book on cryptography by Katz and Lindell (KatzLindell ), Chapter 5.1. Using brute-force search, one must try at most $2^{m}$ different inputs to find a pre-image of a given output. The following corollary of Definition˜6.2 stems from this fact and states that given a prefix of the pre-image, our search space decreases.
Corollary 6.3 (Tunable pre-image runtime).
Let $H_{m}:\{0,1\}^{m}\to\{0,1\}^{m}$ be a random oracle (Definition˜6.2). Fix $k<m$ and a prefix $p\in\{0,1\}^{k}$, and let $H_{m}(p\|r)=y$ be the output of $H_{m}$ for some extension $r$ of $P$. For any (possibly randomized) algorithm $\mathcal{X}$ that makes at most $q$ oracle queries to $H_{m}$ and outputs $z\in\{0,1\}^{m-k}$, it holds that
$\Pr_{H_{m},\mathcal{X}}[H_{m}(p\|z)=y]\leq\frac{q}{2^{m-k}}$
Definition 6.4 (Threshold obfuscation scheme).
Let $\Pi=(Prep,Probe)$ be a pair of algorithms. If given a representation of a threshold function $f_{T}$ (6.1), $Prep$ produces output $O_{T}=Prep(f_{T})$ such that, for any given input $x$ of $f_{T}$, $Probe(O_{T},x)=f_{T}(x)$ holds with high probability, then $\Pi$ is a threshold obfuscation scheme.
We move one to formulate our security definition. We base this definition of the adversarial experiment format used in the book by Katz &\& Lindell KatzLindell . There exists a definition for obfuscation in DBLP:journals/jacm/BarakGIRSVY12 (Boaz Barak et al.). This definition requires that no probabilistic polynomial-time adversary can generate the same outputs as the original program with high probability. While we would like to have such obfuscation for the threshold function (a topic for possible future work), our current obfuscation construction does not satisfy this very strong notion; rather, we require that to reproduce the same outputs, an adversary must have a runtime that is at least logarithmic in the size of the construction, and an adversary must reproduce the exact outputs with probability $1$; otherwise, it cannot do any better than a random guess. Specifically, since we tailor this obfuscation to a threshold function, we formulate this specific notion of obfuscation as an experiment where the adversary gets the obfuscated threshold function just like the processes, and must recover the threshold $T$ in order to succeed.
Definition 6.5 (Obfuscation Security Experiment $Obf_{\mathcal{A},\Pi,\ell}$).
Let $\ell\in\mathbb{N}$ be the range of values for threshold $T$. Let $\Pi=(Prep,Probe)$ be an obfuscation scheme (6.4). Let $\mathcal{T}_{Probe}$ be the runtime of algorithm $Probe$. Let $\mathcal{A}$ be an adversarial algorithm with runtime $\mathcal{T_{A}}<\mathcal{T}_{Probe}\cdot\log\ell$.
  1. 1.
    A random $T\in[\ell]$ is chosen.
  2. 2.
    Compute $O_{T}=Prep(f_{T})$.
  3. 3.
    $\mathcal{A}$ is given $O_{T}$.
  4. 4.
    $\mathcal{A}$ returns an integer $T^{\prime}$.
  5. 5.
    Let $\tau$ be the number of times $\mathcal{A}$ computes $H_{m}$.
  6. 6.
    The output of the experiment is $1,\tau$, if $T^{\prime}=T$, otherwise the output is $0,\tau$.
Using Definition˜6.5, we formulate a definition of security for a threshold obfuscation scheme. While Katz&Lindell would define the security as an upper bound on the probability of success in the experiment $Obf$, in our case, the adversary must be able to cause the protocol to terminate with certainty. We therefore require that the number of hashes the adversary must compute in order to find $T$ with certainty is bounded in expectation.
Definition 6.6.
A threshold obfuscation scheme (Definition˜6.4) $\Pi=(Prep,Probe)$ is secure if for every adversary algorithm $\mathcal{A}$, it holds that
$\mathbb{E}\left[\tau\mid b=1\right]\leq\frac{1}{2}\left((2^{m-k}+1)\log(\ell+1)\right)$
where $b,\tau=Obf_{\mathcal{A},\Pi,\ell}$.
While a bound on the probability of success would be stronger, this definition is sufficient to show that the adversary must do more work than the processes in order to cause the protocol to fail. We move on to define our threshold obfuscation scheme. We begin with the $Prep$ algorithm, the output of which will be used to compute the output of the threshold function in the $Probe$ algorithm. We call this algorithm preprocess.
preprocess 1. Fix integers $1<k<m$ and $\ell\geq 2$. 2. Choose threshold $1<T<\ell$ uniformly at random. 3. For all $1\leq i\leq\ell$, sample binary strings $P_{i}\leftarrow\{0,1\}^{k}$ and $r_{i}\leftarrow\{0,1\}^{m-k}$ uniformly at random. 4. For all $1\leq i\leq\ell$ (a) compute $d=H(P_{i}\|r_{i})$ (b) set $C_{i}\leftarrow d_{1}d_{2}\dots d_{m-1}$ that is, $C_{i}$ is the first m1m-1 bits of $D$ (c) set $V_{i}\leftarrow d_{m}\oplus f_{T}(i)$ 5. Return $(P_{1},\dots,P_{\ell},C_{1},\dots,C_{\ell},v_{1},\dots,v_{\ell})$.
Next, we define the $Probe$ algorithm, which we will use to compute a single output of $f_{T}$.
probe 1. Assume access to $H_{m}$ as well as $C_{1},\dots,C_{\ell}$, $P_{1},\dots,P_{\ell}$, and $v_{1},\dots,v_{\ell}$. 2. Let $1\leq i\leq\ell$ be an input of $f_{T}$. 3. Find string $r\in\{0,1\}^{m-k}$ such that $H_{m}(P_{i}\|r)]_{j}=[C_{i}]_{j}$ for all $j=1,\dots,m-1$. 4. Return $H_{m}(P_{i}\|r)]_{m}\oplus v_{i}$
A more complete definition and implementation can be found in Section˜6.3.
6.2 Correctness &\& security
We prove that Algorithms 6.1 and 6.2 (labeled as preprocess and probe) constitute a threshold obfuscation scheme (6.4), and also satisfy the security definition 6.6. In ˜6.8 we show the correctness of the construction, and in ˜6.11 we show that it is secure. We begin with the following lemma.
Lemma 6.7.
[Distinct pre-image with prefix] Let $k,m\in\mathbb{N}$ be two integers where $k<m$, and let $H_{m}$ be a random oracle 6.2. Let $p\in\{0,1\}^{k},\quad r\in\{0,1\}^{m-k}$ be two binary strings. It holds that
$\Pr\left[\exists_{x\neq r}(H_{m}(p||x)=H_{m}(p||r))\right]\leq 2^{-k}$
Proof.
Let $p\in\{0,1\}^{k},\quad r\in\{0,1\}^{m-k}$ be two binary strings. Choose uniformly at random $x\in\{0,1\}^{m-k}\setminus\{r\}$. By Definition˜6.2, it holds that
$\Pr[H_{m}(p\|r)=H_{m}(p\|x)]\leq 2^{-m}$
By the union bound
$\Pr[\exists x(H_{m}(p\|r)=H_{m}(p\|x))]=\Pr[\bigcup_{x\neq r}H_{m}(p\|r)=H_{m}(p\|x)]$
$\leq\sum_{x\neq r}\Pr[H_{m}(p\|r)=H_{m}(p\|x)]$
Since there are $2^{m-k}-1$ possible choices for $x$ and assuming each sample of $x$ happens independently of the others,
$(2^{m-k}-1)\cdot\Pr[H_{m}(p\|r)=H_{m}(p\|x)]=(2^{m-k}-1)\cdot 2^{-m}\leq 2^{-k}$
Hence, we obtain that
$\Pr[\exists x(H_{m}(p\|r)=H_{m}(p\|x))]\leq 2^{-k}$
The following is the construction correctness claim. We require that the parameters $k,\ell$ satisfy $\frac{\ell}{2^{k}}=negl(k)$, meaning that their ratio is negligible in $k$. This assumption is reasonable since increasing $k$ decreases the probability of a collision as well as reduces the amount of work required to find the suffix of a pre-image.
Claim 6.8.
The algorithm pair $(\textnormal{{preprocess}},\textnormal{{probe}})$ is a threshold obfuscation scheme 6.4, for $\ell,k$ that satisfy $\frac{\ell}{2^{k}}=negl(k)$.
Proof.
Let $\ell\in\mathbb{N}$ be some positive integer and let $f_{T}$ be some threshold function where $0<T<\ell$. Fix some $m,k$ such that $0<k<m$. Let $C,P=\textnormal{{preprocess}}(\ell,T,m,k)$ be the outputs of algorithm preprocess. For each possible input $i\in[\ell]$ of $f_{T}$, the algorithm executes the following steps:
In lines 6.1, 6.1, the algorithm picks random strings $P[i]\in\{0,1\}^{k}$ and $r_{i}\in\{0,1\}^{m-k}$. Their concatenation $P[i]\|r_{i}$ is a string of length $m$ which is a valid input for $H_{m}$.
In 6.1, the algorithm computes the hash of $d\leftarrow H_{m}(P[i]\|r_{i})$.
In 6.1, the algorithm stores the first m1m-1 bits of $d$ in $C[i]$.
In 6.1, the algorithm stores $f_{T}(i)$ in $v_{i}$ by assigning $v_{i}\leftarrow d_{m}\oplus f_{T}(i)$.
Therefore, for any input $i$ of $f_{T}$, there exists at least one string $r_{i}\in\{0,1\}^{m-k}$ such that $H_{m}(P[i]\|r_{i})[j]=C[i][j]$ for all $j=1,\dots,m-1$ and $H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)=V[i]$.
Algorithm probe has access to $\ell,m,k,C,P,V$ and is given input $i$ of the threshold function. The algorithm finds suffix $y$ such that $H_{m}(P[i]\|y)[1,\dots,m-1]=C[i][1,\dots,m-1]$. It remains to show that $H_{m}(P[i]\|y)[m]\oplus V[i]=f_{T}(i)$ holds with high probability. Assume the contrary.
This implies $H_{m}(P[i]\|y)[m]\oplus H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)\neq f_{T}(i)$ which can only happen if $H_{m}(P[i]\|y)[m]\neq H_{m}(P[i]\|r_{i})[m]$. Since $H_{m}$ is deterministic, this implies that $y\neq r_{i}$. But by Lemma˜6.7, the probability that this occurs is at most $2^{-k}$. Thus, for a large enough $k$, it holds that $y=r_{i}$ with high probability. Thus, $H_{m}(P[i]\|y)=H_{m}(P[i]\|r_{i})[m]$, which means that, w.h.p
$H_{m}(P[i]\|y)\oplus V[i]=H_{m}(P[i]\|r_{i})[m]\oplus H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)=f_{T}(i)$
Lemma 6.9.
For any adversary $\mathcal{A}$ that recovers $T$ from $C,P,V$, it holds that
$\mathbb{E}[\tau]\geq\frac{1}{2}\left((2^{m-k}+1)\cdot\log(\ell+1)\right)$
where $\tau$ is the number of times $\mathcal{A}$ computes $H_{m}$.
Proof.
By Definition˜6.2, given $C,P,V$, the threshold $T$ that is hidden in $C,P,V$ can be any value in $\{1,\dots,\ell\}$ with equal probability. Moreover, Definition˜6.2 implies that for each $i\in[\ell]$, every suffix $y\in\{0,1\}^{m-k}$ has equal probability of satisfying $H_{m}(P[i]\|y)[1,\dots,m-1]=C[i]$. It follows that any adversarial algorithm that attempts to compute $T$ directly from $C,P,V$ must check $H_{m}(P[i]\|y)[1,\dots,m-1]=C[i]$ for some number of strings $y$. Order the set $\{0,1\}^{m-k}$ by some arbitrary order (by the number each string represents for example). Denote $X_{i,j}=H_{m}(P[i]\|y)$ where $y$ is the $j^{th}$ string in $\{0,1\}^{m-k}$. Denote $r=2^{m-k}$. Let $M\in\{0,1,\bot,?\}^{\ell\times r}$ be the matrix such that
[…]
Notice that:
  • (a)
    Since the adversary only needs to find one suffix $y$ per column $i$ that satisfies $X_{i,j}[1,\dots,m-1]=C[i]$, it holds that $M_{i,j}=1\vee 0\implies\forall_{g\neq j}M_{i,g}=\bot$.
  • (b)
    Since $C,P,V$ encode the truth table of a threshold function, the adversary can know that if $M_{i,j}=0$ then for all $1\leq r\leq i$ it holds that $M_{r,j^{\prime}}=0$ for some $j^{\prime}$, meaning $f_{T}(r)=0$.
We call the process of checking if a single suffix $y$ satisfies $X_{i,j}[1,\dots,m-1]=C[i]$ a probe of $M$ (not to be confused with the probeprobe algorithm). Denote $M^{(t)}$ as the state of $M$ after the adversarial algorithm makes $t$ probes. Conversely, denote $M^{(0)}$ as the initial state of $M$ at the start of the execution, this implies $M^{(0)}_{i,j}$=? for all $i,j$. Let $L_{t}$ be the number of columns in $M^{(t)}$ with at least one ?? cell (conversely, without any $1$ or 0 cells). Let $P_{t}$ be the number of non-?? cells, in columns with at least one ?? in $M^{(t)}$. Define potential function $\Phi$ as follows
$\Phi(L,P)=\frac{1}{2}\left((r+1)\cdot\log(L+1)-P\right)$
We show that $\Phi$ is the lower bound on the expected number of probes $\mathcal{A}$ must execute to find $T$.
Claim 6.10.
Denote $\text{hit}_{t}$ as the event where the $t^{th}$ probe $\mathcal{A}$ makes is at index $i,j$ and it holds that $X_{i,j}[1,\dots,m-1]=C[i]\wedge(X_{i,j}[m]\oplus V[i]=1\vee X_{i,j}[m]\oplus V[i]=0)$. Conditioned on $\text{hit}_{t}$, it holds that $\log(L_{t+1}+1)\geq\log(L_{t}+1)-1$, that is, a successful probe reduces the number ofpossibilities for $T$ by at most half.
Proof. Since $\mathcal{A}$ knows that $x=\max\{i\mid f_{T}(i)=0\}\leq T\leq\min\{i\mid f_{T}(i)=1\}$, it follows that $T$ is uniform on $\{x,x+1,\dots,x+L_{t}\}$, it follows that
$\Pr[f_{T}(x+j)=0]=\Pr[T>x+j]=\frac{j}{L_{t}+1},\qquad\Pr[f_{T}(x+j)=1]=\Pr[T\leq x+j]=\frac{L_{t}+1-j}{L_{t}+1}$
If $f_{T}(x+j)=0$ then $T\in\{x,\dots,x+j-1\}$ and the new interval size is $j$.
If $f_{T}(x+j)=1$ then $T\in\{x+j,\dots,x+L_{t}\}$ and the new interval size is $L_{t}+1-j$.
This implies that
$\mathbb{E}\left[\log(L_{t+1}+1)\mid\text{hit}_{t}\right]=\frac{j}{L_{t}+1}\log j+\frac{L_{t}+1-j}{L_{t}+1}\log(L_{t}+1-j)$
this expression is minizied at $j=\frac{L_{t}+1}{2}$, thus
$\mathbb{E}\left[\log(L_{t+1}+1)\mid\text{hit}_{t}\right]\geq\log((L_{t}+1)/2)=\log(L_{t}+1)-1$
$\mathbb{E}\left[\log(L_{t+1}+1)-\log(L_{t}+1)\mid\text{hit}_{t}\right]\geq-1\Diamond$
[…]
$\Pr[\text{hit}_{t}]=\frac{1}{r-q_{t}}$
If $\text{hit}_{t}$ does not occur, then $P_{t+1}=P_{t}+1$.
If $\text{hit}_{t}$ does occur, then by (b), we have that $P_{t+1}\leq P_{t}+1-(q_{t}+1)=P_{t}-q_{t}$.
Denote $\Delta P=P_{t+1}-P_{t}$, it follows that
$\mathbb{E}[\Delta P]\leq(1-\Pr[\text{hit}_{t}])+\Pr[\text{hit}_{t}]\cdot(-q_{t})=1-\Pr[\text{hit}_{t}]\cdot(q_{t}+1)$
$\mathbb{E}[\Delta P]\geq\Pr[\text{hit}_{t}]\cdot(q_{t}+1)-1\quad\text{(c)}$
Similarly, denote $\Delta\log(L+1)=\log(L_{t+1}+1)-\log(L_{t}+1)$ and $\Delta\Phi=\Phi(L_{t+1},P_{t+1})-\Phi(L_{t},P_{t})$. It follows that
$\mathbb{E}[\Delta\Phi]=\frac{1}{2}\left((+1)\mathbb{E}[\Delta\log(L+1)]-\mathbb{E}[\Delta P]\right)$
$\frac{1}{2}\left((+1)\Pr[\text{hit}_{t}]\cdot\mathbb{E}[\Delta\log(L+1)\mid\text{hit}_{t}]-\mathbb{E}[\Delta P]\right)$
By ˜6.10 and (c),
$\mathbb{E}[\Delta\Phi]\geq\frac{1}{2}\left((+1)\cdot\Pr[\text{hit}_{t}]\cdot(-1)-1+\Pr[\text{hit}_{t}]\cdot(q_{t}+1)\right)$
$\frac{1}{2}\cdot\left(-1+\Pr[\text{hit}_{t}](q_{t}-r)\right)$
$\frac{1}{2}\cdot\left(-1+\frac{1}{r-q_{t}}(q_{t}-r)\right)$
Thus
$\mathbb{E}[\Delta\Phi]\geq\frac{1}{2}(-1+(-1))=-1\quad\text{(d)}$
Having bounded the expected change in $\Phi$, we bound the expected number of steps $t$ that $\mathcal{A}$ will need to make in order to find $T$. We begin by showing that $\Phi+t$ is submartingale. Notice that when $L_{t}=0$ then the adversary knows all outputs of $f_{T}(\cdot)$, thus we define stopping time as
$\tau=\min\{t\mid L_{t}=0\}$
By definition, if $L_{\tau}=0$ then there are no unknown columns at time $\tau$; thus, $P_{\tau}=0$, which implies that
$\Phi(L_{\tau},P_{\tau})=\Phi(0,0)=\frac{1}{2}\left((+1)\log(0+1)-0\right)=0$
Let $\mathcal{F}_{t}=\sigma(M^{(t)})$ be the filtration generated by $M$ after $t$ probes. Denote $\Phi_{t}=\Phi(L_{t},P_{t})$. By (d), it holds that
$\mathbb{E}[\Phi_{t+1}-\Phi_{t}\mid\mathcal{F}_{t}]\geq-1$
$\implies\mathbb{E}[\Phi_{t+1}-\Phi_{t}+1\mid\mathcal{F}_{t}]\geq 0$
$\implies\mathbb{E}[(\Phi_{t+1}+t)-(\Phi_{t}+t)+1\mid\mathcal{F}_{t}]\geq 0$
$\implies\mathbb{E}[(\Phi_{t+1}+(t+1))-(\Phi_{t}+t)\mid\mathcal{F}_{t}]\geq 0$
Since $\Phi_{t}$ is completely determined by $M^{(t)}$, $\mathbb{E}[\Phi_{t}\mid\mathcal{F}_{t}]=\Phi_{t}$. This implies that
$\mathbb{E}[(\Phi_{t+1}+(t+1))-(\Phi_{t}+t)\mid\mathcal{F}_{t}]=\mathbb{E}[(\Phi_{t+1}+(t+1))\mid\mathcal{F}_{t}]-\mathbb{E}[(\Phi_{t}+t)\mid\mathcal{F}_{t}]\geq 0$
$\implies\mathbb{E}[(\Phi_{t+1}+(t+1))\mid\mathcal{F}_{t}]\geq\mathbb{E}[(\Phi_{t}+t)\mid\mathcal{F}_{t}]=\Phi_{t}+t\quad(e)$
Therefore, $\Phi_{t}+t$ is submartingale. Notice that for $\ell\cdot r$ probes, the adversary knows all columns of $M$ and the algorithm stops, thus it follows that $\tau\leq\ell\cdot r<\infty$ is bounded. Therefore, the optional stopping theorem applies to $\Phi_{t}+t$. By (e), and the optional stopping theorem
$\mathbb{E}[\Phi_{\tau}+\tau]\geq\Phi_{0}$
since $\Phi_{\tau}=0$
$\implies\mathbb{E}[\tau]\geq\Phi_{0}=\frac{1}{2}\left((2^{m-k}+1)\cdot\log(\ell+1)\right)$
The following is a security claim based on the established definition.
Claim 6.11.
The threshold function obfuscation $\Pi=\textnormal{{preprocess}},\textnormal{{probe}}$ is secure 6.6.
Proof.
Let $\mathcal{A}$ be any adversarial algorithm that tries to recover $T$ from $C,P,V$. Define $\tau,b=Obf_{\mathcal{A},\Pi,\ell}$ By Lemma˜6.9, if $\mathcal{A}$ finds $T$ and does $\tau$ probes to $H_{m}$, it holds that
$\mathbb{E}[\tau]\leq\frac{1}{2}\left((2^{m-k}+1)\log(\ell+1)\right)$
therefore
$\mathbb{E}[\tau\mid b=1]\leq\frac{1}{2}\left((2^{m-k}+1)\log(\ell+1)\right)$
6.3 Implementation
Here we give a more detailed description of the obfuscated threshold implementation outlined in Section 6.1.
6.3.1 Outline
We present an implementation of the first algorithm of our threshold obfuscation scheme, Algorithm 6.1. This algorithm accepts hash size parameter $m$, hardness parameter $k$, value range $\ell$, and input threshold $T$. It returns three arrays, $C$, $P$ and $V$. The idea is to hide the threshold $T$ in an array of bits $V=v_{1},\dots,v_{\ell}$. We create an array of hashes $C=C_{1},\dots,C_{\ell}$ where each hash is created by randomly choosing two strings $p_{i}\in\{0,1\}^{k}$ and $r_{i}\in\{0,1\}^{m-k}$ for each $i=1,\dots,\ell$ and hashing their concatenation $p_{i}\|r_{i}$. We store in $C_{i}$ the first m1m-1 bits of the hash $H_{m}(p_{i}\|r_{i})$ and the last bit we XOR with the value of $f_{T}(i)$ to obscure it. We store this in the bit $v_{i}\leftarrow[H_{m}(p_{i}\|r_{i})]_{m}\oplus f_{T}(i)$. In this way, in order to compute $f_{T}(i)$, one must find a string $y$ such that $H_{m}(p_{i}\|y)]_{1\dots(m-1)}=C_{i}$, and then then compute $H_{m}(p_{i}\|y)]_{m}\oplus v_{i}$. Given that there is a high probability that $H_{m}(p_{i}\|y)]_{1,\dots,(m-1)}=[H_{m}(p_{i}\|r_{i})]_{1,\dots,(m-1)}\implies[H_{m}(p_{i}\|y)]_{m}=[H_{m}(p_{i}\|r_{i})]_{m}$ it follows that, $H_{m}(p_{i}\|y)]_{m}\oplus v_{i}=[H_{m}(p_{i}\|y)]_{m}\oplus[H_{m}(p_{i}\|r_{i})]_{m}\oplus f_{T}(i)=f_{T}(i)$ holds w.h.p. To create the hashes, we employ our random oracle function (6.2). Since one needs to find a suffix for a pre-image of the hash (up to the last bit) in order find $f_{T}(i)$, we must ensure that finding such suffix is feasible. To control the size of the search space we have a security parameter $k$ which determines the size of the prefix string $p_{i}$. For this reason, the preprocessing algorithm returns the array of prefix strings $P=P_{1},\dots,P_{\ell}$. Increasing $k$ decreases the size of the suffix required to find, which automatically reduces the search space. By the properties of a random oracle (6.2), reducing the size of the search space implies reducing the runtime required to find $r_{i}$.
In a practical implementation of the construction, the standard hash pub2015secure SHA 512/256 can be used in place of $H_{m}$ (Definition˜6.2). While no formal proof exists, it is widely accepted that in practice, SHA behaves like a random oracle, including an avalanche effect, where any minor change in the input, even a single bit, drastically alters the output.
1procedure $\textnormal{{preprocess}}(\ell,T,m,k$ s.t. $t\leq\ell\wedge k<m)$
 // Initialize array of hashes and array of nonces
2 $C[1,...,\ell][1,...m-1]$ an uninitizalized array of bit strings
3 $P[1,...,\ell][1,...k]$ an uninitizalized array of bit strings
[…]
5 for $i\textleftarrow 1\dots\ell$ do
6    $P[i]\leftarrow$ random bit string of length $k$
7    $r_{i}\leftarrow$ random bit string of length mkm-k
8    $d\leftarrow H_{m}(P[i]\|r_{i}$)
9    $C[i]\leftarrow d_{1}d_{2}\dots d_{m-1}$
10    $V[i]\leftarrow d_{m}\oplus f_{T}(i)$
11 return $C,P,V$
12
Algorithm 6.1 Threshold Encapsulation
6.3.2 Preprocessing Algorithm
Algorithm 6.1 is executed by the programmer to obtain an obfuscated protocol. The integer parameters $\ell$, $T$, $k$, and $m$ are declared (Line 1). Where $\ell$ is the span in which a threshold $t$ is defined, $m$ is the number of bits used by the one-way function, say 256 for SHA-256, and $k$ is the length of the (exposed) prefix nonce of the $m$ bits input (and output) of the SHA.
Two arrays, $C$ (committed values) and $P$ (padding nonces), each of $\ell$ entries, are defined in lines 6.1 and 6.1. Each entry of the $C$ array is ready to be assigned with the SHA result (line 6.1). Each entry of the $P$ array exposes a $k$-bit nonce value (line 6.1). We use the nonce to enforce the independence of pre-image searches and as randomness used to validate the pre-image.
The for loop (lines 6.1 to 6.1) iterates over all possible values of $i$ and computes the cryptographic hash (e.g., the common one-way function such as SHA) over the nonce concatenated with $r_{i}$, a concealed random string of length mkm-k. The computation starts (line 6.1) with a selection of random bit string of size $k$ as the nonce for $i$ and entry and stores it in $P[i]$. Next, the algorithm randomly selects a bit string $r_{i}$ (line 6.1); Then, the algorithm computes the hash with the concatenation of the two strings as input (line 6.1). Then stores the first m1m-1 bits of the hash in $C[i]$ (line 6.1), and stores the XOR of the last bit with $f_{T}(i)$ in $V[i]$ (line 6.1). By the time the algorithm reaches the return command (line 6.1), it has computed the hashes for all the $\ell$ hash values of $f_{T}$, and all the nonces used to calculate them are stored in $P$; thus, the algorithm returns them to be embedded in the protocol program.
local data: integers $\ell,m,k$, precomputed array of hashes $C[1,...,\ell][1,...,m-1]$, prefixes $P[1,...,\ell][1,...,k]$, and bits $V[1,\dots,\ell]$
1 procedure $\textnormal{{probe}}(i\in[\ell])$
2 Y[1,,mk]Y[1,...,m-k] an uninitialized bit string
3 D[1,,m]D[1,...,m] an uninitialized bit string
4 $Y\leftarrow$ string that satisfies $C[i][1,\dots,m-1]=H_{m}(P[i]\|Y)$
5 $D\leftarrow H_{m}(P[i]\|Y)$
6 return $D[m]\oplus V[i]$
7
Algorithm 6.2 Threshold Probe
6.3.3 Probing algorithm
The processes use the probe procedure described in Algorithm 6.2 to compute $f_{T}(i)$.
The algorithm works by finding the pre-image $P[i]\|r_{i}$ that was used to compute $C[i]$ in the preprocessing phase. It begins by initializing a bit string $Y$ of length mkm-k to hold potential suffixes to pre-image $P[i]$ of $C[i]$, and an $m$-length bit string $D$ to hold the hash that is equal to $C[i]$ in the first m1m-1 bits. Next, it finds and stores in $Y$ a suffix for the pre-image of $C[i]$ such that the first m1m-1 bits of the hash of $P[i]\|Y$ are equal to $C[i]$ (line 6.2). It stores this hash in $D$ (line 6.2). This means that, with high probability $D[m]=H_{m}(P[i]\|r_{i})[m]$. The exact method of finding this pre-image has not been specified, but exhaustive search mining or any other method to find a pre-image of a cryptographic hash function can be employed. Thus, Algorithm 6.2 returns $f_{T}(i)$ by computing $D[m]\oplus V[i]$ (line 6.2), since the preprocessing algorithm defined $V[i]=H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)$, the algorithm returns $f_{T}(i)$.
7 Conclusion
Designing an algorithm to cope with an adversarial scheduler (and/or inputs) ensures the algorithm’s function in rare scenarios where the schedule is the most unfortunate for the algorithm. In many cases, the worst case leads to an impossibility result, while the rare scenario does not happen in practice. This is particularly true in scenarios that require ongoing tracing and computation of the imaginary scheduler entity.
We examine the power of using program obfuscations and random oracles to derandomize distributed asynchronous consensus algorithms. The power of the random oracle can be demonstrated in algorithms where symmetry is broken by other means, such as process identifiers and/or symmetry-breaking operations, such as compare-and-swap. Thus, randomization has an inherent more substantial power in breaking symmetry. On the other hand, harvesting proper randomization during runtime is challenging, and it should be avoided if possible.
A weakness of our construction is that it requires substantial preprocessing to construct an obfuscated threshold function since our implementation is just a truth table hidden behind time-lock puzzles. A natural question is whether a more sophisticated obfuscation procedure could reduce this setup, perhaps by constructing a function $f(s,i)$ where the threshold $T$ is a hash of some shared data $s$ but the implementation of $f$ prevents recovering $T$ more efficiently than simply doing binary search.
However, even with these limitations, we believe that we have demonstrated integrating program obfuscation and random oracle abstractions and functionalities in distributed computing is helpful for the theory and practice of distributed computing and systems.